### v1

# CodeGRAG: Extracting Composed Syntax Graphs for Retrieval Augmented Cross-Lingual Code Generation

Abstract
Utilizing large language models to generate codes has shown promising meaning in software development revolution. Despite the intelligence shown by the general large language models, their specificity in code generation can still be improved due to the syntactic gap and mismatched vocabulary existing among natural language and different programming languages. In addition, programming languages are inherently logical and complex, making them hard to be correctly generated. Existing methods rely on multiple prompts to the large language model to explore better solutions, which is expensive. In this paper, we propose Syntax Graph Retrieval Augmented Code Generation (CodeGRAG) to enhance the performance of LLMs in single-round code generation tasks. CodeGRAG extracts and summarizes the control flow and data flow of code blocks to fill the gap between programming languages and natural language. The extracted external structural knowledge models the inherent flows of code blocks, which can facilitate LLMs for better understanding of code syntax and serve as a bridge among different programming languages. CodeGRAG significantly improves the code generation ability of LLMs and can even offer performance gain for cross-lingual code generation, e.g., C++ for Python.
1 Introduction
In recent years, large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023a) have shown great impact in various domains. Automated code generation emerges as a captivating frontier (Zheng et al., 2023; Roziere et al., 2023; Shen et al., 2023), promising to revolutionize software development by enabling machines to write and optimize code with minimal human intervention.
Despite the success in natural language understanding and generation, there is still a gap between the natural language and programming language as illustrated in Figure 1. Code, unlike natural language, possesses a well-defined structure that governs its syntax and semantics. This structure provides valuable information about the relationships between different parts of the code, the flow of execution, and the overall organization of the functions Jiang et al. (2021); Guo et al. (2020). General-purpose LLMs regard a code block as a sequence of tokens. By ignoring the inherent structure of codes, they miss out on essential cues that could help them better understand and generate code. In addition, the multi-lingual code generation abilities of LLMs remain a challenge due to the gap between different programming languages. Moreover, the training data of different programming languages can be unbalanced, which leads to a biased ability of multi-lingual code generation.
Figure 1: Illustration of the gap between the programming language and the natural language.
Previous methods on code representation learning seek to interpret code blocks based on the abstract syntax tree, data flow graph (Aho et al., 2006), control flow graph (Allen, 1970), and the intermediate states (Lattner and Adve, 2004), where the semantic level and logical level information of codes are better captured. These methods gain the performance improvement in code classification (Long et al., 2022), code search (Zeng et al., 2023), etc, showcasing the benefits of capturing the structures of codes. In this paper, we aim to inject the structural knowledge of codes into the general-purpose large language models for a better understanding of codes. By understanding the underlying structure of code, the LLMs can 1) extract the inherent semantic and logic flow of codes for higher quality generation and 2) transfer knowledge across different programming languages, which enables cross-lingual code generation.
However, tuning the base model for structural code understanding is expensive and heavy. Recently, retrieval augmented generation (RAG) has shown great performance gains in facilitating LLMs on complex and temporal-related tasks, well utilizing the in-context-learning ability to instruct LLMs for enhanced generation. In this paper, we propose to develop a code RAG framework to retrieve relevant structural code representations to fill the gap between natural language and programming language, provide hints for enhanced code generation, and utilize the structure information of codes to offer cross-lingual code generation augmentation.
The main challenges of designing an effective framework of RAG for code generation can be summarized in threefolds. 1) How to design effective structural views of code blocks to represent the inherent syntax of codes, which could fill the gap between natural language and programming languages, inform LLMs of external syntactic knowledge, and model the underlying cross-lingual knowledge. 2) How to measure the structural similarities between codes and how to align the natural language and programming language for joint similarity computation. 3) How to obtain an informative query format for efficient knowledge querying, which can offer consistent useful information extraction under different sets of problems that are diversified in syntax and styles.
Concretely, we propose to 1) represent the code blocks with the composed syntax graph based on the data-flow-graph and control-flow-graph, which extracts both the semantic level and the logical level information of code blocks, 2) use hybrid-GNN and pretrained cross-lingual code search models to measure the distances, and 3) obtain the problem description and the function declaration as the query content.
The main contributions of the paper can be summarized as follows:
  • We propose an efficient and effective method named CodeGRAG that bridges the gap among natural language and programming languages, transfers knowledge among different programming languages, and utilizes the in-context learning ability of LLMs for enhanced code generation. The generation process requires only one calling of LLMs and can offer cross-lingual enhancement.
  • We propose an effective graphical view to purify the semantic and logic knowledge from the code space, which offers more useful information than the raw code block and can summarize the cross-lingual knowledge.
  • We propose an effective retrieval method to retrieve relevant code knowledge for generation enhancement, where the cross-modality alignment and structure preserving are achieved.
2 Related Work
Figure 2: The illustration of the overall process of CodeGRAG.
LLMs for NL2Code. The evolution of the Natural Language to Code translation (NL2Code) task has been significantly influenced by the development of large language models (LLMs). Initially, general LLMs like GPT-J (Radford et al., 2023), GPT-NeoX (Black et al., 2022), and LLaMA (Touvron et al., 2023a), despite not being specifically tailored for code generation, showed notable NL2Code capabilities due to their training on datasets containing extensive code data like the Pile (Gao et al., 2020) and ROOTS (Laurençon et al., 2022). To further enhance these capabilities, additional pretraining specifically focused on code has been employed. PaLM-Coder, an adaptation of the PaLM model (Chowdhery et al., 2023), underwent further training on an extra 7.8 billion code tokens, significantly improving its performance in code-related tasks. Similarly, Code LLaMA (Roziere et al., 2023) represents an advancement of LLaMA2 (Touvron et al., 2023b), benefiting from extended training on over 500 billion code tokens, leading to marked improvements over previous models in both code generation and understanding. These developments underscore the potential of adapting generalist LLMs to specific domains like NL2Code through targeted training, leading to more effective and efficient code translation solutions.
Code Search. The code search methods can be summarized into three folds. Early methods utilizes sparse search to match the query and codes (Hill et al., 2011; Yang and Huang, 2017), which suffers from mismatched vocabulary due to the gap between natural language and codes. Neural methods (Cambronero et al., 2019; Gu et al., 2021) then focus on mapping the query and codes into a joint representation space for more accurate retrieval. With the success of pretrained language models, many methods propose to use pretraining tasks to improve the code understanding abilities and align different language spaces. For example, CodeBERT (Feng et al., 2020) is pretrained on NL-PL pairs of 6 programming languages with the masked language modeling and replaced token detection task. CodeT5 (Wang et al., 2021) supports both code-related understanding and generation tasks through bimodal dual generation. UniXcoder (Guo et al., 2022) integrates the aforementioned pretraining tasks, which is a unified cross-modal pre-trained model.
Code Representation. Early methods regard code snippets as sequences of tokens, assuming the adjacent tokens will have strong correlations. This line of methods (Harer et al., 2018; Ben-Nun et al., 2018; Feng et al., 2020; Ciniselli et al., 2021) take programming languages as the same with the natural language, using language models to encode the code snippets too. However, this ignoring of the inherent structure of codes leads to a loss of expressiveness. Methods that take the structural information of codes into consideration then emerge. Mou et al. (2016) used convolution networks over the abstract syntax tree (AST) extracted from codes. Alon et al. (2019) encoded paths sampled from the AST to represent codes. Further exploration into the graphical representation of codes (Allamanis et al., 2017) is conducted to better encode the structures of codes, where more intermediate states of the codes are considered.
3 Methodology
Figure 3: Illustration of the extracted composed syntax graph from the code block. The arrows in the bottom part indicate the names of different edges, which are extracted based on the ASTs.
3.1 Overview
In this paper, we leverage both generative models and retrieval models to produce results that are both coherent and informed by the external syntax knowledge. The overall process of CodeGRAG is illustrated in Figure 2, which consists of three stages as follows.
Syntax Knowledge Base Preparation. We analyze the syntax and control information of code blocks and extract their graphical views to better represent the codes. This process can be formulated as, $\forall c_{i}\in D_{\text{pool}}$:
$g_{i}\xleftarrow{}\text{GraphExtractor}(c_{i}),$ (1)
[…]
where $c_{i}$ is the raw code block and $g_{i}$ is the extracted graphical view of the corresponding code block.
Knowledge Querying. Given a target problem to be completed, we generate informative query of it and use it to retrieve knowledge from the constructed knowledge base. The process can be formulated as:
$q\xleftarrow{}\text{QueryExtractor}(p),$ (3)
[…]
where $q$ denotes the query content, $p$ denotes the target problem, and $i$ is the returned index of the Top-1 relevant content stored in the constructed knowledge base.
Knowledge Augmented Generation. After obtaining the knowledge index, we access the corresponding entry of the knowledge base and obtain the graphical syntax knowledge to inform the LLMs for enhanced code generation. The process can be formulated as:
$c_{i},g_{i}$ $\xleftarrow{}\text{KB}[i],$ (5)
result $\xleftarrow{}\text{LLM}(p,g_{i}).$ (6)
3.2 Syntax Graph Knowledge Base
In this section, we discuss how to extract informative views for code blocks and how to summarize the retrieved knowledge. An illustration is given in Figure 3. To capture both the semantic and the logical information, we propose to combine the data flow graph (Aho et al., 2006) and the control flow graph (Allen, 1970) with the read-write signals (Long et al., 2022) to represent the code blocks, both of them are constructed on the base of the abstract syntax tree.
Abstract Syntax Tree (AST). An abstract syntax tree (AST) is a tree data structure that represents the abstract syntactic structure of source code. An AST is constructed by a parser, which reads the source code and creates a tree of nodes. Each node in the tree represents a syntactic construct in the source code, such as a statement, an expression, or a declaration. ASTs are used in a variety of compiler construction and program analysis tasks, including: parsing, type checking, optimization, and program analysis. ASTs have good compactness and can represent the structure of the source code in a clear and concise way.
Data Flow Graph (DFG). The data flow graph (DFG) is a graphical representation of the flow of data dependencies within a program. It is a directed graph that models how data is transformed and propagated through different parts of a program. In DFG, nodes are operands and edges indicate data flows. Two types of edges are considered: 1) operation edges that connect the nodes to be operated and the nodes that receive the operation results; 2) function edges that indicate data flows for function calls and returns. These edges connect nodes, including non-temporary operands and temporary operands, which refer to variables and constants that explicitly exist in the source code, and variables existing only in execution, respectively.
Control Flow Graph (CFG). The control flow graph (CFG) is a graphical representation of the flow of control or the sequence of execution within a program. It is a directed graph that models the control relationships between different parts of a program. Based on compiler principles, we slightly adjust the design of CFG to better capture the key information of the program. Nodes in CFG are operations in the source code, including standard operations, function calls and returns. Edges indicate the execution order of operations.
Composed Syntax Graph. A composed syntax graph composes the data flow graph and the control flow graph with the read-write flow existing in the code blocks. An illustration of the extracted composed syntax graph is displayed in Figure 3. Different edge types along with their concrete names are given in colors. As for the node names, the middle figure displays the concrete types of nodes (operands) and the right figure displays the properties of nodes.
After obtaining the composed syntax graphs for code blocks in the retrieval pool, we use them for retrieval enhancement, where both the semantic level and the logic level information are preserved.
3.3 Knowledge Querying
The main problems of the retrieval lie in: 1) how to fill the gap between natural languages and programming language and the gap among different programming languages; 2) how to properly measure the similarity of code blocks, e.g., the unique structure existing in programming languages; 3) how to design the informative query content.
3.3.1 Query Content
Since the styles of different code problems can diversify, the query content of the retrieval process matters. We consider the following contents: 1) Problem description, which describes the task to be completed by the target function code. Potential ambiguity and style diversity may exist among different problems set, which lead to a decrease in retrieval accuracy. 2) Function declaration, which gives the function name and the input variables.
Before knowledge querying, we first extract the problem description of each task to reduce the ambiguity and then concatenate it with the function declaration to serve as the query content, where the functionality and input format of the expected code block are contained.
3.3.2 Retrieval Function
The query of the retrieval includes problem description $Q_{p}$ and function description $Q_{c}$, while each content of the retrieval pool includes raw code block $V_{c}$ and its graphical view $V_{g}$.
To expressively represent each of the components, we use the pretrained NL2Code model encoder $\phi_{1}(\cdot)$ to represent the problem description and code snippets and design a graph neural network $\phi_{2}(\cdot)$ to represent the graphical view. The retrieval function can be formulated as:
$\mathbf{h^{V}}$ $=\psi(\phi_{1}(V_{c}),\phi_{2}(V_{g})),$ (7)
$\mathbf{h^{Q}}$ $=\phi_{1}(Q_{p}\|Q_{c}),$ (8)
Distance $=1-\frac{\mathbf{\mathbf{h^{Q}}\cdot\mathbf{h^{V}}}}{\|\mathbf{h^{Q}}\|\cdot\|\mathbf{h^{V}}\|}.$ (9)
We use models pretrained to align natural language and programming language to serve as the encoder $\phi_{1}$:
  • CodeT5+ (Wang et al., 2023) is an advanced encoder-decoder foundation LLM, tailored for diverse code understanding and generation tasks. It is designed for various code-related tasks, capable of operating in different modes (encoder-only, decoder-only, encoder-decoder). It employs a range of pretraining tasks, including span denoising, causal language modeling on code, and text-code contrastive learning, matching and CLM tasks. This approach improves its performance in retrieval tasks by effectively aligning text and code data.
  • UnixCoder (Guo et al., 2022) is a cross-modal pre-trained model optimized for both code understanding and generation tasks. Leveraging a multi-layer Transformer structure with mask attention matrices, it outperforms traditional unified encoder-decoder models in auto-regressive tasks like code completion. UniXcoder uniquely employs multi-modal inputs, including code comments and abstract syntax trees (AST), enhancing its code representation capabilities. It employs three types of language modeling tasks: masked language modeling, unidirectional language modeling, and denoising objectives. Additionally, it introduces two advanced pre-training tasks: multi-modal contrastive learning and cross-modal generation, which refine the semantic representation of code fragments.
And we design a graph neural network $\phi_{2}$ to preserve the semantic and logical information of code blocks. The representation of each node $\mathbf{n}_{i}^{(0)}$ and edge $\mathbf{e}_{i}^{(0)}$ are first initialized with vectors corresponding to the node text and edge text encoded by $\phi_{1}$. A message passing process is first conducted to fuse the semantic and structural information into each node representation.
$\mathbf{m}_{ij}^{(l)}=\mathbf{W}^{(l)}$ $(\mathbf{n}_{i}^{(l-1)}\|\mathbf{e}_{ij}^{(l-1)}),$ (10)
$Q_{j}^{(l)}=$ $\mathbf{W_{Q}}^{(l)}\mathbf{n}_{j}^{(l-1)},$ (11)
$\mathbf{K}_{ij}^{(l)}=\mathbf{W_{K}}^{(l)}m_{ij}^{(l)},$  $\text{\quad}\mathbf{V}_{ij}^{(l)}=\mathbf{W_{V}}^{(l)}\mathbf{m}_{ij}^{(l)},$ (12)
$a_{ij}^{(l)}=\text{softmax}$ (${}_{i\in N(j)}\mathbf{(Q}_{j}^{(l)}\mathbf{K}_{ij}^{(l)}),$ (13)
$\mathbf{n}_{j}^{(l)}=\sum_{i\in N(j)}$ $a_{ij}^{(l)}\mathbf{V}_{ij}^{(l)}.$ (14)
A global attention-based readout is then applied to obtain the graph representation:
$\mathbf{g}=\sum_{i}\text{softmax}(f_{\text{gate}}(\mathbf{n}_{i}^{L}))f_{\text{feat}}(\mathbf{n}_{i}^{L}).$ (15)
The different encoding networks are optimized via the contrastive learning based self-supervised training, which includes the intra-modality contrastive learning and inter-modality contrastive learning. The intra-modality constrastive learning serves for preserving the modality information, while the inter-modality contrastive learning serves for modality alignment.
Alignment Contrastive Learning. There are two types of alignment to be ensured during the retrieval process: 1) Question-Answer (QA) alignment and 2) Code-Graph (CG) alignment. We define the positive pairs for QA alignment purpose as $\mathcal{I}^{+}_{QA}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{i}^{Q}\rangle|i\in D_{\text{train}}\}$ and define the negative pairs for QA alignment purpose as $\mathcal{I}^{-}_{QA}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{j}^{Q}\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$
And we define the positive pairs for CG alignment purpose as $\mathcal{I}^{+}_{CG}=\{\langle\phi_{1}(c_{i}),\phi_{2}(g_{i})\rangle|i\in D_{\text{train}}\}$ and define the negative pairs for CG alignment purpose as $\mathcal{I}^{-}_{CG}=\{\langle\phi_{1}(c_{i}),\phi_{2}(g_{j})\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$
Structure Preserving Contrastive Learning. To preserve the structural information of the graphical views, we perform intra-modality contrastive learning among the graphical views and their corrupted views. Concretely, we corrupt each of the graphical view $g_{i}$ with the edge dropping operation to obtain its corrupted view $g_{i}^{\prime}$. The positive pairs for structure-preserving purpose are then designed as $\mathcal{I}^{+}_{\text{preserve}}=\{\langle\phi_{2}(g_{i}),\phi_{2}(g_{i}^{\prime})\rangle|i\in D_{\text{train}}\}$. The negative pairs for structure preserving purpose are designed as $\mathcal{I}^{-}_{\text{preserve}}=\{\langle\phi_{2}(g_{i}),\phi_{2}(g_{j}^{\prime})\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}$.
3.4 Knowledge Augmented Generation
After we obtain the summarized retrieved knowledge, we append it to the prompt text for knowledge augmented generation. The prompt used is illustrated in Figure 4.
Figure 4: The illustration of the used prompt.
The prompt consists of three components: the system prompt illustrated in the blue part, the retrieved knowledge and hints illustrated in the green part, and the problem (including task description, function declaration, etc.) illustrated in the yellow part. The three parts are concatenated to be fed into LLMs for knowledge augmented generation.
Table 1: Main results of Pass@1 on Humaneval-X.
Model Size Model Humaneval-X C++ Humaneval-X Python
6B GPT-J 7.54 11.10
6B CodeGen-Multi 11.44 19.41
6.7B InCoder 9.50 16.41
13B CodeGeeX 17.06 22.89
16B CodeGen-Multi 18.05 19.22
16B CodeGen-Mono 19.51 29.28
15B StarCoder 31.55 32.93
15B WizardCoder 29.27 57.3
15B Pangu-Coder2 45.12 64.63
- GPT-3.5-Turbo 57.93 71.95
- GPT-3.5-Turbo + Code-RAG 60.37 72.56
- GPT-3.5-Turbo + CodeGRAG 62.20 72.56
- GPT-3.5-Turbo + Cross-Lingual Code-RAG 62.20 70.12
- GPT-3.5-Turbo + Cross-Lingual CodeGRAG 64.02 77.44
4 Experiments
In this section, we conduct experiments to answer the following research questions (RQs).
  • RQ1
    Does the proposed CodeGRAG offer performance gain against the base model?
  • RQ2
    Does the proposed graph view abstract more informative knowledge compared with the raw code block?
  • RQ3
    Does the proposed CodeGRAG model the high-level thought-of-codes? Can CodeGRAG offer cross-lingual augmentation?
  • RQ4
    How does the retrieval algorithms impact the performance?
  • RQ5
    What is the impact of each of the components of the graphical view?
  • RQ6
    How is the compatibility of CodeGRAG?
4.1 Setup
In this paper, we evaluate CodeGRAG with the widely used HumanEval-X (Zheng et al., 2023) dataset, which is a multi-lingual code benchmark. We use greedy decoding strategy to do the generation. The evaluation metric is Pass@1.
We evaluate the multi-lingual code generation abilities of 1) models with less than 10 billion parameters: GPT-J (Radford et al., 2023), CodeGen-Multi (Nijkamp et al., 2022), and InCoder(Fried et al., 2022); 2) models with 10-20 billion parameters: CodeGeeX (Zheng et al., 2023), CodeGen-Multi (Nijkamp et al., 2022), CodeGen-Mono (Nijkamp et al., 2022), StarCoder (Li et al., 2023), WizardCoder (Luo et al., 2023), and Pangu-Coder2 (Shen et al., 2023); 3) close-sourced GPT-3.5 model. The detailed statistics of the model sizes are displayed in Table 1.
4.2 Main Results
The main results are summarized in Table 1. Code-RAG represents directly using the raw code block as the retrieved knowledge and CodeGRAG refers to using the graphical view as the knowledge. From the results, we can draw the following conclusions:
RQ1. The proposed CodeGRAG could offer performance gain against the base model, which validates the effectiveness of the proposed RAG for code generation framework.
RQ2. The model with CodeGRAG could beat model with raw code block augmented generation. From the results, we can see that the proposed graph view could summarize the useful structural syntax information and filter out the noises, which could offer more informative knowledge hints than the raw code block.
RQ3. Python code generation is improved with cross-lingual retrieval, which uses C++ code and corresponding graphical views as the retrieved knowledge for an enhanced generation. Therefore, the proposed CodeGRAG could offer cross-lingual augmentation. This validates that the proposed graphical view could abstract the high-level and inherent information (e.g., the control and data flow to solve a specific problem) of the code blocks, which are shared across different programming languages regarding solving the same problem.
Table 2: Results with different content encoders (yielding different retrieval models).
Content Encoder $\phi_{1}(\cdot)$ Method Humaneval-X C++ Humaneval-X Python
NA GPT-3.5-Turbo 57.93 71.95
CodeT5 + Code-RAG 60.37 72.56
+ CodeGRAG 62.20 72.56
+ Cross-Lingual Code-RAG 61.59 70.12
+ Cross-Lingual CodeGRAG 62.20 76.83
UnixCoder + Code-RAG 58.54 71.95
+ CodeGRAG 62.20 71.95
+ Cross-Lingual Code-RAG 62.20 66.46
+ Cross-Lingual CodeGRAG 64.02 77.44
4.3 Impact of Retrieval Methods (RQ4)
In this section, we explore how different retrieval methods influence the final performance. We use CodeT5 and UnixCoder as the text and code encoder, which are both trained with multi-lingual data and dedicated for multi-lingual search. The results are summarized in Table 2.
From the results, we could see that the augmentation performance can differ with different retrieval content encoders. Specifically, the cross-lingual augmentation performance increases with UnixCoder being the encoder compared with that of CodeT5 being encoder. Since UnixCoder involves in the abstract syntax tree for structure capturing, while CodeT5 does not. This validates the effectiveness of involving structure information. In addition, we can see that with better cross-lingual retrieval accuracy, directly using raw codes as cross-lingual hints will bias the model generation, while feeding graphical views of syntax will boost the model performance.
4.4 Impacts of the Components of the Graphical View (RQ5)
In this section, we adjust the inputs of the graph components to the LLMs as the external knowledge. Concretely, we study the information contained in node names, edge names, and the topological structure. The results are presented in Table 3.
Table 3: The impacts of the graph components.
Datasets Python C++
Edge Type Only 73.78 61.59
Edge Type + Node Name 75.00 59.76
Edge Type + Node Type 75.61 59.15
Edge Type + Topological 77.44 64.02
The edge type refers to the type of flows between operands (child, read, write, etc.), the node type refers to the type of operands (DeclStmt, temp, etc.), the node name refers to the name of the intermediate variables, and the topological information refers to the statistics of the concrete numbers of different types of edges. From the results, we can observe that 1) the edge features matter the most in constructing the structural view of code blocks for enhancement, 2) the type of nodes expresses the most in representing operands information, and 3) the overall structure of the graphical view also gives additional information.
4.5 Discussion of the Compatibility (RQ6)
Despite the effectiveness of the proposed graphical views to represent the code blocks, the flexibility and convenience of applying the graphical views extraction process is important for wider application of the proposed method. In this section, we discuss the compatibility of CodeGRAG.
First of all, the extraction process of all the graphical views are front-end. Therefore, this extraction process applies to a wide range of code, even error code. One could also use convenient tools to reformulate the code and improve the pass rate of the extraction process.
In addition, we give the ratio of generated results that can pass the graphical views extraction process, which is denoted by Extraction Rate. The Pass@1 and the Extraction Rate of the generated results passing the graphical extraction process are given in Table 4 and Table 5.
Table 4: The extraction rate of the generated results passing the graphical extraction process (C++).
Generated Codes Pass@1 Extraction Rate
Code-RAG 62.20 92.07
CodeGRAG 64.02 92.68
Table 5: The extraction rate of the generated results passing the graphical extraction process (Python).
Generated Codes Pass@1 Extraction Rate
Code-RAG 71.95 91.46
CodeGRAG 77.44 96.95
From the results, we could see that the extraction rates are high for codes to pass the graphical views extraction process, even under the situation where the Pass@1 ratios of the generated results are low. This indicates that the application range of the proposed method is wide. In addition, as the code RAG also offers performance gains, one could use multiple views, including raw code blocks and graphical representations, as the retrieval knowledge base.
5 Conclusion
Despite the expanding role of LLMs in code generation, there are inherent challenges pertaining to their understanding of code syntax and their multi-lingual code generation capabilities. This paper introduces the Syntax Graph Retrieval Augmented Code Generation (CodeGRAG) to enhance LLMs for single round and cross-lingual code generation. CodeGRAG extracts and summarizes data flow and control flow information from codes, effectively bridging the gap between programming language and natural language. By integrating external structural knowledge, CodeGRAG enhances LLMs’ comprehension of code syntax and empowers them to generate complex and multi-lingual code with improved accuracy and fluency. The method’s demonstrated performance gains signal its potential to significantly elevate LLMs’ code generation abilities, including facilitating cross-lingual programming language generation such as C++ to Python. Overall, the findings warrant optimism regarding the transformative impact of CodeGRAG in advancing LLMs’ proficiency in code generation across various programming languages.
Limitations
In this paper, we propose a tuning-free method that can offer enhanced single round and cross-lingual code generation. Despite the efficiency and effectiveness, there are also limitations within this work. Firstly, dependency on the quality of the external knowledge base could be a potential concern. The quality of the external knowledge base could be improved with regular expression extraction on the noisy texts and codes. Secondly, the potential of combining the RAG with finetuning for improved generation is not explored in this work. Finetuning the LLMs with the assistance of CodeGRAG could be a future exploration.
Appendix A Example of the inserted graphical view
An illustration of the inserted graphical view is given below.
{mdframed}
Graph(
num_nodes=’node’: 24,
num_edges=(’node’, ’-0’, ’node’): 1, (’node’, ’-1’, ’node’): 1, (’node’, ’ArraySubscriptExpredge0’, ’node’): 1, (’node’, ’ArraySubscriptExpredge1’, ’node’): 1, (’node’, ’CXXOperatorCallExpredge1’, ’node’): 1, (’node’, ’CXXOperatorCallExpredge2’, ’node’): 2, (’node’, ’ImplicitCastExpredge0’, ’node’): 1, (’node’, ’UserDefineFun’, ’node’): 1, (’node’, ’falseNext’, ’node’): 1, (’node’, ’next’, ’node’): 5, (’node’, ’read’, ’node’): 10, (’node’, ’trueNext’, ’node’): 1, (’node’, ’write’, ’node’): 9,
metagraph=[(’node’, ’node’, ’-0’), (’node’, ’node’, ’-1’), (’node’, ’node’, ’ArraySubscriptExpredge0’), (’node’, ’node’, ’ArraySubscriptExpredge1’), (’node’, ’node’, ’CXXOperatorCallExpredge1’), (’node’, ’node’, ’CXXOperatorCallExpredge2’), (’node’, ’node’, ’ImplicitCastExpredge0’), (’node’, ’node’, ’UserDefineFun’), (’node’, ’node’, ’falseNext’), (’node’, ’node’, ’next’), (’node’, ’node’, ’read’), (’node’, ’node’, ’trueNext’), (’node’, ’node’, ’write’)]

### v2

@@
-CodeGRAG: Extracting Composed Syntax Graphs for Retrieval Augmented Cross-Lingual Code Generation
+CodeGRAG: Bridging the Gap between Natural Language and Programming Language via Graphical Retrieval Augmented Generation
 Abstract
@@
 Despite the intelligence shown by the general large language models, their specificity in code generation can still be improved due to the syntactic gap and mismatched vocabulary existing among natural language and different programming languages.
-In addition, programming languages are inherently logical and complex, making them hard to be correctly generated.
-Existing methods rely on multiple prompts to the large language model to explore better solutions, which is expensive.
-In this paper, we propose Syntax Graph Retrieval Augmented Code Generation (CodeGRAG) to enhance the performance of LLMs in single-round code generation tasks.
-CodeGRAG extracts and summarizes the control flow and data flow of code blocks to fill the gap between programming languages and natural language.
-The extracted external structural knowledge models the inherent flows of code blocks, which can facilitate LLMs for better understanding of code syntax and serve as a bridge among different programming languages.
-CodeGRAG significantly improves the code generation ability of LLMs and can even offer performance gain for cross-lingual code generation, e.g., C++ for Python.
+In this paper, we propose CodeGRAG, a Graphical Retrieval Augmented Code Generation framework to enhance the performance of LLMs.
+CodeGRAG builds the graphical view of code blocks based on the control flow and data flow of them to fill the gap between programming languages and natural language, which can facilitate natural language based LLMs for better understanding of code syntax and serve as a bridge among different programming languages.
+To take the extracted structural knowledge into the foundation models, we propose 1) a hard meta-graph prompt template to transform the challenging graphical representation into informative knowledge for tuning-free models and 2) a soft prompting technique that injects the domain knowledge of programming languages into the model parameters via finetuning the models with the help of a pretrained GNN expert model.
+Various experiments and ablations are done on four datasets including both the C++ and python languages to validate the hard meta-graph prompt, the soft prompting technique, and the effectiveness of the objectives for pretrained GNN expert.
+CodeGRAG improves the code generation ability of LLMs and can even offer performance gain for cross-lingual code generation.
+The implementation is available at https://anonymous.4open.science/r/Code-5970/ .
 1 Introduction
@@
 Automated code generation emerges as a captivating frontier (Zheng et al., 2023; Roziere et al., 2023; Shen et al., 2023), promising to revolutionize software development by enabling machines to write and optimize code with minimal human intervention.
-Despite the success in natural language understanding and generation, there is still a gap between the natural language and programming language as illustrated in Figure 1.
+However, syntactic gap and mismatched vocabulary among natural language (NL) and programming languages (PL) exist, hindering LLM’s performance on code generation.
+As illustrated in Figure 1, programming language (marked in blue) contains special tokens such as “int” or “++” that natural language (marked in yellow) doesn’t possess, leading to vocabulary mismatch.
+Besides, the relations between tokens in programming languages are often structural, e.g., the complex branching and jumps, whereas natural language is arranged simply in sequential manner, leading to syntactic gap.
+For example, in the control flow graph of the raw code (marked in pink), two “if” blocks (marked in purple) are adjacent and are executed sequentially under certain condition, but they appear to be intervaled in raw textual code.
+As discussed above, the innate structures of programming languages are different from that of the sequential-based natural language.
+The challenges of enhancing a general-purposed large language models for code-related tasks can be summarized into two folds.
+Figure 1: Illustration of the gap between the programming language and the natural language.
+(C1) How to solve the gap between different languages and better interpret the inherent logic of code blocks.
 Code, unlike natural language, possesses a well-defined structure that governs its syntax and semantics.
@@
 By ignoring the inherent structure of codes, they miss out on essential cues that could help them better understand and generate code.
-In addition, the multi-lingual code generation abilities of LLMs remain a challenge due to the gap between different programming languages.
-Moreover, the training data of different programming languages can be unbalanced, which leads to a biased ability of multi-lingual code generation.
-Figure 1: Illustration of the gap between the programming language and the natural language.
-Previous methods on code representation learning seek to interpret code blocks based on the abstract syntax tree, data flow graph (Aho et al., 2006), control flow graph (Allen, 1970), and the intermediate states (Lattner and Adve, 2004), where the semantic level and logical level information of codes are better captured.
-These methods gain the performance improvement in code classification (Long et al., 2022), code search (Zeng et al., 2023), etc, showcasing the benefits of capturing the structures of codes.
-In this paper, we aim to inject the structural knowledge of codes into the general-purpose large language models for a better understanding of codes.
-By understanding the underlying structure of code, the LLMs can 1) extract the inherent semantic and logic flow of codes for higher quality generation and 2) transfer knowledge across different programming languages, which enables cross-lingual code generation.
-However, tuning the base model for structural code understanding is expensive and heavy.
-Recently, retrieval augmented generation (RAG) has shown great performance gains in facilitating LLMs on complex and temporal-related tasks, well utilizing the in-context-learning ability to instruct LLMs for enhanced generation.
-In this paper, we propose to develop a code RAG framework to retrieve relevant structural code representations to fill the gap between natural language and programming language, provide hints for enhanced code generation, and utilize the structure information of codes to offer cross-lingual code generation augmentation.
-The main challenges of designing an effective framework of RAG for code generation can be summarized in threefolds.
-1) How to design effective structural views of code blocks to represent the inherent syntax of codes, which could fill the gap between natural language and programming languages, inform LLMs of external syntactic knowledge, and model the underlying cross-lingual knowledge.
-2) How to measure the structural similarities between codes and how to align the natural language and programming language for joint similarity computation.
-3) How to obtain an informative query format for efficient knowledge querying, which can offer consistent useful information extraction under different sets of problems that are diversified in syntax and styles.
-Concretely, we propose to 1) represent the code blocks with the composed syntax graph based on the data-flow-graph and control-flow-graph, which extracts both the semantic level and the logical level information of code blocks, 2) use hybrid-GNN and pretrained cross-lingual code search models to measure the distances, and 3) obtain the problem description and the function declaration as the query content.
+In addition, the multi-lingual code generation abilities of LLMs is challenging due to the gap among different programming languages.
+(C2) How to inject the innate knowledge of programming languages into general purpose large language models for enhancement.
+Despite the well representation of the programming knowledge, the ways to inject the knowledge into the NL-based foundation models is also challenging.
+The structural representation of codes could be hard to understand, which poses a challenge to the capability of the foundation models.
+To solve the above challenges, we propose CodeGRAG, a graphical retrieval augmented generation framework for code generation.
+For (C1), we propose to interpret the code blocks using the composed graph based on the data-flow and control-flow of the code block, which extracts both the semantic level and the logical level information of the code.
+The composed graphical view could 1) better capture the innate structural knowledge of codes for NL-based language models to understand and 2) model the innate function of code blocks that bridging different programming languages.
+For (C2), we propose a meta-graph prompting technique for tuning-free models and a soft-prompting technique for tuned models.
+The meta-graph prompt summarizes the overall information of the extracted graphical view and transforms the challenging and noisy graphical representation into informative knowledge.
+The soft-prompting technique deals with the graphical view of codes with a pretrained GNN expert network and inject the processed knowledge embedding into the parameters of the general-purpose foundation models with the help of supervised finetuning.
 The main contributions of the paper can be summarized as follows:
-  • We propose an efficient and effective method named CodeGRAG that bridges the gap among natural language and programming languages, transfers knowledge among different programming languages, and utilizes the in-context learning ability of LLMs for enhanced code generation.
-The generation process requires only one calling of LLMs and can offer cross-lingual enhancement.
+  • Novel GraphRAG framework for code generation.
+We propose CodeGRAG that bridges the gap among natural language and programming languages, transfers knowledge among different programming languages, and enhances the ability of LLMs for code generation.
+CodeGRAG requires only one calling of LLMs and can offer multi-lingual enhancement.
+  • Effective graphical view to inform and stimulate the structural programming knowledge of LLMs.
 We propose an effective graphical view to purify the semantic and logic knowledge from the code space, which offers more useful information than the raw code block and can summarize the cross-lingual knowledge.
-  • We propose an effective retrieval method to retrieve relevant code knowledge for generation enhancement, where the cross-modality alignment and structure preserving are achieved.
-2 Related Work
-Figure 2: The illustration of the overall process of CodeGRAG.
-LLMs for NL2Code.
-The evolution of the Natural Language to Code translation (NL2Code) task has been significantly influenced by the development of large language models (LLMs).
-Initially, general LLMs like GPT-J (Radford et al., 2023), GPT-NeoX (Black et al., 2022), and LLaMA (Touvron et al., 2023a), despite not being specifically tailored for code generation, showed notable NL2Code capabilities due to their training on datasets containing extensive code data like the Pile (Gao et al., 2020) and ROOTS (Laurençon et al., 2022).
-To further enhance these capabilities, additional pretraining specifically focused on code has been employed.
-PaLM-Coder, an adaptation of the PaLM model (Chowdhery et al., 2023), underwent further training on an extra 7.8 billion code tokens, significantly improving its performance in code-related tasks.
-Similarly, Code LLaMA (Roziere et al., 2023) represents an advancement of LLaMA2 (Touvron et al., 2023b), benefiting from extended training on over 500 billion code tokens, leading to marked improvements over previous models in both code generation and understanding.
-These developments underscore the potential of adapting generalist LLMs to specific domains like NL2Code through targeted training, leading to more effective and efficient code translation solutions.
-Code Search.
-The code search methods can be summarized into three folds.
-Early methods utilizes sparse search to match the query and codes (Hill et al., 2011; Yang and Huang, 2017), which suffers from mismatched vocabulary due to the gap between natural language and codes.
-Neural methods (Cambronero et al., 2019; Gu et al., 2021) then focus on mapping the query and codes into a joint representation space for more accurate retrieval.
-With the success of pretrained language models, many methods propose to use pretraining tasks to improve the code understanding abilities and align different language spaces.
-For example, CodeBERT (Feng et al., 2020) is pretrained on NL-PL pairs of 6 programming languages with the masked language modeling and replaced token detection task.
-CodeT5 (Wang et al., 2021) supports both code-related understanding and generation tasks through bimodal dual generation.
-UniXcoder (Guo et al., 2022) integrates the aforementioned pretraining tasks, which is a unified cross-modal pre-trained model.
-Code Representation.
-Early methods regard code snippets as sequences of tokens, assuming the adjacent tokens will have strong correlations.
-This line of methods (Harer et al., 2018; Ben-Nun et al., 2018; Feng et al., 2020; Ciniselli et al., 2021) take programming languages as the same with the natural language, using language models to encode the code snippets too.
-However, this ignoring of the inherent structure of codes leads to a loss of expressiveness.
-Methods that take the structural information of codes into consideration then emerge.
-Mou et al.
-(2016) used convolution networks over the abstract syntax tree (AST) extracted from codes.
-Alon et al.
-(2019) encoded paths sampled from the AST to represent codes.
-Further exploration into the graphical representation of codes (Allamanis et al., 2017) is conducted to better encode the structures of codes, where more intermediate states of the codes are considered.
+  • Effective soft prompting technique to preserve the programming domain knowledge and inject it into LLMs parameters.
+We propose an effective soft prompting technique, which injects the domain knowledge of programming languages into the model parameters via finetuning LLMs with the assistance of a pretrained GNN expert model.
 2 Methodology
-Figure 3: Illustration of the extracted composed syntax graph from the code block. The arrows in the bottom part indicate the names of different edges, which are extracted based on the ASTs.
+Figure 2: Overview of CodeGRAG.
+(Top) Knowledge Preparation.
+We extract the control flow and data flow of each external code block and compose them using the read-write signal to obtain the semantic and logical expression of each code block, which is then abstracted into graphical view as hard knowledge document and embedded into GraphEmb as soft knowledge document.
+The GraphEmb is encoded by a pretrained GNN expert model constrained by the alignment and structure preserving objectives.
+(Bottem) Retrieval Augmented Generation.
+We extract query from the task input and retrieve from the external corpus.
+For tuning free models, we use the hard graphical view to stimulate the structural programming knowledge of LLMs for enhanced generation.
+For tunable models, we use the soft GraphEmb and inject the programming domain knowledge into LLMs parameters via finetuning them with the GNN expert signals.
+The expert signals informed LLMs can then produce enhanced generation.
 2.1 Overview
-In this paper, we leverage both generative models and retrieval models to produce results that are both coherent and informed by the external syntax knowledge.
-The overall process of CodeGRAG is illustrated in Figure 2, which consists of three stages as follows.
-Syntax Knowledge Base Preparation.
+In this paper, we leverage both generative models and retrieval models to produce results that are both coherent and informed by the expert graphical knowledge of programming language.
+The overall process of CodeGRAG is illustrated in Figure 2, which mainly consists of three stages: graphical knowledge base preparation, knowledge querying, and graphical knowledge augmented generation.
+2.2 Graphical Knowledge Base Preparation
+In this section, we discuss how to extract informative graphical views for code blocks.
 We analyze the syntax and control information of code blocks and extract their graphical views to better represent the codes.
@@
 […]
-where $c_{i}$ is the raw code block and $g_{i}$ is the extracted graphical view of the corresponding code block.
-Knowledge Querying. Given a target problem to be completed, we generate informative query of it and use it to retrieve knowledge from the constructed knowledge base. The process can be formulated as:
-$q\xleftarrow{}\text{QueryExtractor}(p),$ (3)
-[…]
-where $q$ denotes the query content, $p$ denotes the target problem, and $i$ is the returned index of the Top-1 relevant content stored in the constructed knowledge base.
-Knowledge Augmented Generation.
-After obtaining the knowledge index, we access the corresponding entry of the knowledge base and obtain the graphical syntax knowledge to inform the LLMs for enhanced code generation.
-The process can be formulated as:
-$c_{i},g_{i}$ $\xleftarrow{}\text{KB}[i],$ (5)
-result $\xleftarrow{}\text{LLM}(p,g_{i}).$ (6)
-3.2 Syntax Graph Knowledge Base
-In this section, we discuss how to extract informative views for code blocks and how to summarize the retrieved knowledge.
-An illustration is given in Figure 3.
+where $c_{i}$ is the raw code block and $g_{i}$ is the corresponding extracted graphical view.
 To capture both the semantic and the logical information, we propose to combine the data flow graph (Aho et al., 2006) and the control flow graph (Allen, 1970) with the read-write signals (Long et al., 2022) to represent the code blocks, both of them are constructed on the base of the abstract syntax tree.
@@
 These edges connect nodes, including non-temporary operands and temporary operands, which refer to variables and constants that explicitly exist in the source code, and variables existing only in execution, respectively.
+Figure 3: Illustration of the extracted composed syntax graph from the code block. The arrows in the bottom part indicate the names of different edges, which are extracted based on the ASTs.
 Control Flow Graph (CFG).
@@
 As for the node names, the middle figure displays the concrete types of nodes (operands) and the right figure displays the properties of nodes.
-After obtaining the composed syntax graphs for code blocks in the retrieval pool, we use them for retrieval enhancement, where both the semantic level and the logic level information are preserved.
+An illustration of the composed graphical view is given in Figure 3.
+After obtaining the composed syntax graphs for code blocks in the retrieval pool, we use them to inform the general-purpose LLMs to bridge the gap between NL and PLs, where both the semantic level and the logic level information are preserved.
 2.3 Knowledge Querying
-The main problems of the retrieval lie in: 1) how to fill the gap between natural languages and programming language and the gap among different programming languages; 2) how to properly measure the similarity of code blocks, e.g., the unique structure existing in programming languages; 3) how to design the informative query content.
-3.3.1 Query Content
+Given a target problem to be completed, we generate informative query of it and use it to retrieve graphical knowledge from the constructed knowledge base. The process can be formulated as:
+$q\xleftarrow{}\text{QueryExtractor}(p),$ (3)
+[…]
+where $q$ denotes the query content, $p$ denotes the target problem, and $i$ is the returned index of the Top-1 relevant content stored in the constructed knowledge base.
+The main problems of the retrieval lie in: 1) how to design the informative query content and 2) how to align the different modalities.
+2.3.1 Query Extractor
 Since the styles of different code problems can diversify, the query content of the retrieval process matters.
@@
 Before knowledge querying, we first extract the problem description of each task to reduce the ambiguity and then concatenate it with the function declaration to serve as the query content, where the functionality and input format of the expected code block are contained.
-3.3.2 Retrieval Function
+2.3.2 Retriever
 The query of the retrieval includes problem description $Q_{p}$ and function description $Q_{c}$, while each content of the retrieval pool includes raw code block $V_{c}$ and its graphical view $V_{g}$.
-To expressively represent each of the components, we use the pretrained NL2Code model encoder $\phi_{1}(\cdot)$ to represent the problem description and code snippets and design a graph neural network $\phi_{2}(\cdot)$ to represent the graphical view.
-The retrieval function can be formulated as:
-$\mathbf{h^{V}}$ $=\psi(\phi_{1}(V_{c}),\phi_{2}(V_{g})),$ (7)
-$\mathbf{h^{Q}}$ $=\phi_{1}(Q_{p}\|Q_{c}),$ (8)
+To expressively represent the components, we use the encoder $\phi(\cdot)$ of the pretrained NL2Code model to represent the problem description and code snippets. The retrieval function is:
+$\mathbf{h^{V}}$ $=\phi(V_{c}\|V_{g}),$ (5)
+$\mathbf{h^{Q}}$ $=\phi(Q_{p}\|Q_{c}),$ (6)
 Distance $=1-\frac{\mathbf{\mathbf{h^{Q}}\cdot\mathbf{h^{V}}}}{\|\mathbf{h^{Q}}\|\cdot\|\mathbf{h^{V}}\|}.$ (7)
-We use models pretrained to align natural language and programming language to serve as the encoder $\phi_{1}$:
-  • CodeT5+ (Wang et al., 2023) is an advanced encoder-decoder foundation LLM, tailored for diverse code understanding and generation tasks.
-It is designed for various code-related tasks, capable of operating in different modes (encoder-only, decoder-only, encoder-decoder).
-It employs a range of pretraining tasks, including span denoising, causal language modeling on code, and text-code contrastive learning, matching and CLM tasks.
-This approach improves its performance in retrieval tasks by effectively aligning text and code data.
-  • UnixCoder (Guo et al., 2022) is a cross-modal pre-trained model optimized for both code understanding and generation tasks.
-Leveraging a multi-layer Transformer structure with mask attention matrices, it outperforms traditional unified encoder-decoder models in auto-regressive tasks like code completion.
-UniXcoder uniquely employs multi-modal inputs, including code comments and abstract syntax trees (AST), enhancing its code representation capabilities.
-It employs three types of language modeling tasks: masked language modeling, unidirectional language modeling, and denoising objectives.
-Additionally, it introduces two advanced pre-training tasks: multi-modal contrastive learning and cross-modal generation, which refine the semantic representation of code fragments.
-And we design a graph neural network $\phi_{2}$ to preserve the semantic and logical information of code blocks.
+2.4 Graphical Knowledge Augmented Generation
+After we obtain the returned graphical view, we inject it to the foundation LLMs for graphical knowledge augmented generation.
+Since the graphical view is hard to understand, we propose 1) a meta-graph template to transform the graphical view into informative knowledge for tuning-free model and 2) a soft prompting technique to tune the foundation models for their better understanding of the graphical views with the assistance of an expert GNN model.
+2.4.1 Hard Meta-Graph Prompt
+The original graphical view of a code block could contain hundreds of nodes and edges.
+A full description of it could cost a overly long context, along with the understanding challenge posed by the long edge lists.
+Therefore, we propose to use a meta-graph template to abstract the information of the graphical view.
+The abstracted meta-graph consists of the canonical edge types and node types, which describes the basic topology of the graphical view (Sun & Han, 2013), with the textual features obtained from the ASTs contained in the node and edge features.
+Then we use the meta-graph template to transform the retrieved graphical view into digestable knowledge and insert it into the final prompt for generation.
+As illustrated in Figure 5, the final prompt consists of three components: the system prompt illustrated in the blue part, the retrieved knowledge and hints illustrated in the green part, and the problem (including task description, function declaration, etc.) illustrated in the yellow part.
+The three parts are concatenated to be fed into LLMs for knowledge augmented generation.
+Figure 4: Hard meta-graph prompt.
+Figure 5: Soft prompting.
+2.4.2 Soft Prompting with the Expert
+Directly hard prompt to the LLMs poses a challenge to the digesting capability of the backbone LLMs, which could fail under the case where the backbone LLMs cannot well understand the graph components.
+To compress the graphical knowledge into model parameters and help the backbone LLMs to better understand the programming language, we propose a soft prompting technique.
+The overall procedure can summarized into expert encoding of graphical views, finetuning with the expert signal, and inference.
+Expert Encoding of Graphical Views.
+We design a graph neural network to preserve the semantic and logical information of code blocks.
 The representation of each node $\mathbf{n}_{i}^{(0)}$ and edge $\mathbf{e}_{i}^{(0)}$ are first initialized with vectors corresponding to the node text and edge text encoded by $\phi_{1}$.
@@
 $\mathbf{g}=\sum_{i}\text{softmax}(f_{\text{gate}}(\mathbf{n}_{i}^{L}))f_{\text{feat}}(\mathbf{n}_{i}^{L}).$ (13)
-The different encoding networks are optimized via the contrastive learning based self-supervised training, which includes the intra-modality contrastive learning and inter-modality contrastive learning.
+The expert encoding network is optimized via the contrastive learning based self-supervised training, which includes the intra-modality contrastive learning and inter-modality contrastive learning.
 The intra-modality constrastive learning serves for preserving the modality information, while the inter-modality contrastive learning serves for modality alignment.
   • Alignment Contrastive Learning.
-There are two types of alignment to be ensured during the retrieval process: 1) Question-Answer (QA) alignment and 2) Code-Graph (CG) alignment.
-We define the positive pairs for QA alignment purpose as $\mathcal{I}^{+}_{QA}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{i}^{Q}\rangle|i\in D_{\text{train}}\}$ and define the negative pairs for QA alignment purpose as $\mathcal{I}^{-}_{QA}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{j}^{Q}\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$
+There are two types of alignment to be ensured: 1) NL-Code (NC) alignment and 2) Code-Graph (CG) alignment.
+We define the positive pairs for NC alignment purpose as $\mathcal{I}^{+}_{NC}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{i}^{Q}\rangle|i\in D_{\text{train}}\}$ and define the negative pairs for NC alignment purpose as $\mathcal{I}^{-}_{NC}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{j}^{Q}\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$
     And we define the positive pairs for CG alignment purpose as $\mathcal{I}^{+}_{CG}=\{\langle\phi_{1}(c_{i}),\phi_{2}(g_{i})\rangle|i\in D_{\text{train}}\}$ and define the negative pairs for CG alignment purpose as $\mathcal{I}^{-}_{CG}=\{\langle\phi_{1}(c_{i}),\phi_{2}(g_{j})\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$
@@
 The negative pairs for structure preserving purpose are designed as $\mathcal{I}^{-}_{\text{preserve}}=\{\langle\phi_{2}(g_{i}),\phi_{2}(g_{j}^{\prime})\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}$.
-3.4 Knowledge Augmented Generation
-After we obtain the summarized retrieved knowledge, we append it to the prompt text for knowledge augmented generation. The prompt used is illustrated in Figure 4.
-Figure 4: The illustration of the used prompt.
-The prompt consists of three components: the system prompt illustrated in the blue part, the retrieved knowledge and hints illustrated in the green part, and the problem (including task description, function declaration, etc.) illustrated in the yellow part.
-The three parts are concatenated to be fed into LLMs for knowledge augmented generation.
-Table 1: Main results of Pass@1 on Humaneval-X.
-Model Size Model Humaneval-X C++ Humaneval-X Python
-6B GPT-J 7.54 11.10
-6B CodeGen-Multi 11.44 19.41
-6.7B InCoder 9.50 16.41
-13B CodeGeeX 17.06 22.89
-16B CodeGen-Multi 18.05 19.22
-16B CodeGen-Mono 19.51 29.28
-15B StarCoder 31.55 32.93
-15B WizardCoder 29.27 57.3
-15B Pangu-Coder2 45.12 64.63
-- GPT-3.5-Turbo 57.93 71.95
-- GPT-3.5-Turbo + Code-RAG 60.37 72.56
-- GPT-3.5-Turbo + CodeGRAG 62.20 72.56
-- GPT-3.5-Turbo + Cross-Lingual Code-RAG 62.20 70.12
-- GPT-3.5-Turbo + Cross-Lingual CodeGRAG 64.02 77.44
+Finetuning with the Expert Soft Signal.
+To help the backbone LLMs to digest the graphical views, we tune the LLMs with the expert soft signal using supervised finetuning.
+The prompt for finetuning consists of the system prompt, retrieved knowledge where the expert encoded graphical view is contained using a token embedding, and task prompt, which is illustrated in Figure 5.
+Inference. After the finetuning stage, we used the tuned models to generate codes using the soft prompting template as illustrated in Figure 5.
 3 Experiments
-In this section, we conduct experiments to answer the following research questions (RQs).
   • RQ1
@@
   • RQ3
-    Does the proposed CodeGRAG model the high-level thought-of-codes? Can CodeGRAG offer cross-lingual augmentation?
+    Can soft prompting enhance the capability of the backbone LLMs? Does finetuning with the soft prompting outperforms the simple supervised finetuning?
   • RQ4
-    How does the retrieval algorithms impact the performance?
+    Are the proposed pretraining objectives for the GNN expert effective?
   • RQ5
@@
   • RQ6
-    How is the compatibility of CodeGRAG?
+    How is the compatibility of the graphical view?
 3.1 Setup
-In this paper, we evaluate CodeGRAG with the widely used HumanEval-X (Zheng et al., 2023) dataset, which is a multi-lingual code benchmark.
+In this paper, we evaluate CodeGRAG with the widely used HumanEval-X (Zheng et al., 2023) dataset, which is a multi-lingual code benchmark and CodeForce dataset in which we collect real-world programming problems from codeforces [^1] https://codeforces.com/ website.
+For CodeForce dataset we include problems categorized by different difficulty levels corresponding to the website and select $469$ problems of difficulty level A for testing.
 We use greedy decoding strategy to do the generation.
 The evaluation metric is Pass@1.
-We evaluate the multi-lingual code generation abilities of 1) models with less than 10 billion parameters: GPT-J (Radford et al., 2023), CodeGen-Multi (Nijkamp et al., 2022), and InCoder(Fried et al., 2022); 2) models with 10-20 billion parameters: CodeGeeX (Zheng et al., 2023), CodeGen-Multi (Nijkamp et al., 2022), CodeGen-Mono (Nijkamp et al., 2022), StarCoder (Li et al., 2023), WizardCoder (Luo et al., 2023), and Pangu-Coder2 (Shen et al., 2023); 3) close-sourced GPT-3.5 model.
-The detailed statistics of the model sizes are displayed in Table 1.
+For the size of retrieval pool, we use 11,913 C++ code snippets and 2,359 python code snippets.
+Due to the limited access, we do not use a large retrieval corpus for our experiment, which can be enlarged by other people for better performance.
+We also attach the graph extraction codes for both languages and all other expeirment codes here: https://anonymous.4open.science/r/Code-5970/
+For the fintuning details, the learning rate and weight decay for the expert GNN training is 0.001 and 1e-5, repectively.
+We apply 8-bit quantization and use LoRA for parameter-efficient fine-tuning.
+The rank of the low-rank matrices in LoRA is uniformly set to 8, alpha set to 16, and dropout is set to 0.05.
+The LoRA modules are uniformly applied to the Q and V parameter matrices of the attention modules in each layer of the LLM.
+All the three models are optimized using the AdamW optimizer.
+For the CodeContest dataset, totally 10609 datapoints are used, and for APPS dataset, 8691 data samples are used to train the model.
 3.2 Main Results
-The main results are summarized in Table 1.
-Code-RAG represents directly using the raw code block as the retrieved knowledge and CodeGRAG refers to using the graphical view as the knowledge.
+The main results are summarized in Table 1 and Table 2.
+Table 1: Results of Hard Meta-Graph Prompt on Humaneval-X. (Pass@1)
+Model Retrieved Knowledge C++ Python
+GPT-3.5-Turbo N/A 57.93 71.95
+Code Block (Nashid et al., 2023; Lu et al., 2022) 60.37 72.56
+Meta-Graph 62.20 72.56
+(Multi-Lingual) Code-Block (Nashid et al., 2023; Lu et al., 2022) 62.20 70.12
+(Multi-Lingual) Meta-Graph 64.02 77.44
+GPT-4omini N/A 65.24 78.66
+Code Block (Nashid et al., 2023; Lu et al., 2022) 65.24 78.66
+Meta-Graph 65.85 79.88
+(Multi-Lingual) Code-Block (Nashid et al., 2023; Lu et al., 2022) 65.85 79.27
+(Multi-Lingual) Meta-Graph 67.07 80.49
+Table 2: Results of Soft Prompting. (Pass@1)
+Model Finetune CodeForce (C++) APPS (Python)
+Gemma 7b N/A 12.83 5.09
+SFT 14.76 21.09
+Soft Prompting 19.13 26.15
+Llama2 13b N/A 9.61 7.29
+SFT 11.88 12.06
+Soft Prompting 13.62 12.74
+CodeLlama 7b N/A 5.20 24.41
+SFT 9.87 26.15
+Soft Prompting 11.09 30.26
 From the results, we can draw the following conclusions:
-RQ1. The proposed CodeGRAG could offer performance gain against the base model, which validates the effectiveness of the proposed RAG for code generation framework.
+RQ1.
+The proposed CodeGRAG could offer performance gain against the base model, which validates the effectiveness of the proposed graphical retrieval augmented generation for code generation framework.
 RQ2.
-The model with CodeGRAG could beat model with raw code block augmented generation.
+The model informed by the meta-graph (CodeGRAG) could beat model informed by the raw code block.
 From the results, we can see that the proposed graph view could summarize the useful structural syntax information and filter out the noises, which could offer more informative knowledge hints than the raw code block.
+In addition, inserting the intermediate representations of codes into the prompt can stimulate the corresponding programming knowledge of LLMs.
 RQ3.
-Python code generation is improved with cross-lingual retrieval, which uses C++ code and corresponding graphical views as the retrieved knowledge for an enhanced generation.
-Therefore, the proposed CodeGRAG could offer cross-lingual augmentation.
-This validates that the proposed graphical view could abstract the high-level and inherent information (e.g., the control and data flow to solve a specific problem) of the code blocks, which are shared across different programming languages regarding solving the same problem.
-Table 2: Results with different content encoders (yielding different retrieval models).
-Content Encoder $\phi_{1}(\cdot)$ Method Humaneval-X C++ Humaneval-X Python
-NA GPT-3.5-Turbo 57.93 71.95
-CodeT5 + Code-RAG 60.37 72.56
-+ CodeGRAG 62.20 72.56
-+ Cross-Lingual Code-RAG 61.59 70.12
-+ Cross-Lingual CodeGRAG 62.20 76.83
-UnixCoder + Code-RAG 58.54 71.95
-+ CodeGRAG 62.20 71.95
-+ Cross-Lingual Code-RAG 62.20 66.46
-+ Cross-Lingual CodeGRAG 64.02 77.44
-4.3 Impact of Retrieval Methods (RQ4)
-In this section, we explore how different retrieval methods influence the final performance.
-We use CodeT5 and UnixCoder as the text and code encoder, which are both trained with multi-lingual data and dedicated for multi-lingual search.
-The results are summarized in Table 2.
-From the results, we could see that the augmentation performance can differ with different retrieval content encoders.
-Specifically, the cross-lingual augmentation performance increases with UnixCoder being the encoder compared with that of CodeT5 being encoder.
-Since UnixCoder involves in the abstract syntax tree for structure capturing, while CodeT5 does not.
-This validates the effectiveness of involving structure information.
-In addition, we can see that with better cross-lingual retrieval accuracy, directly using raw codes as cross-lingual hints will bias the model generation, while feeding graphical views of syntax will boost the model performance.
+From Table 2, we can see that finetuning with the expert soft prompting could offer more performance gain than that brought by simple supervised finetuning.
+This validates the effectiveness of the designed pretraining expert network and the technique of finetuning with soft prompting, which injects the programming domain knowledge into the LLMs parameters and inform the models with the structural information for gap filling.
+3.3 Impacts of the pretraining objectives for the expert GNN (RQ4)
+To study the effectiveness of the proposed pretraining objectives for the expert GNN, we remove each objective to yield different expert GNNs. The results are in Table 3.
+Table 3: Ablation studies on the GNN pretraining losses.
+Model Finetune CodeForce (C++) APPS (Python)
+Gemma 7b Soft Prompting 19.13 26.15
+w/o Alignment 7.88 28.58
+w/o Structure-Preserving 11.70 21.50
+Llama2 13b Soft Prompting 13.62 12.74
+w/o Alignment 11.79 10.76
+w/o Structure-Preserving 5.50 11.09
+CodeLlama 7b Soft Prompting 11.09 30.26
+w/o Alignment 10.92 29.45
+w/o Structure-Preserving 10.66 26.59
+From the results, we could see that both the Alignment and the Structure Preserving contribute to the expressiveness of the expert GNN model.
+The alignment pretraining objective helps to promote the alignment among natural language, programming language, and their graphical views.
+The structure preserving objective helps to preserve the innate data-flows and control-flows information of code blocks.
+The two objectives collaborate with each other to yield expressive programming domain knowledge GNN expert model, which encodes external programming knowledge and injects the knowledge into LLMs parameters.
 3.4 Impacts of the Components of the Graphical View (RQ5)
-In this section, we adjust the inputs of the graph components to the LLMs as the external knowledge.
+In this section, we adjust the inputs of the graphical components to the LLMs.
 Concretely, we study the information contained in node names, edge names, and the topological structure.
@@
 From the results, we can observe that 1) the edge features matter the most in constructing the structural view of code blocks for enhancement, 2) the type of nodes expresses the most in representing operands information, and 3) the overall structure of the graphical view also gives additional information.
-4.5 Discussion of the Compatibility (RQ6)
+3.5 Compatibility Discussion of the Graphical Views(RQ6)
 Despite the effectiveness of the proposed graphical views to represent the code blocks, the flexibility and convenience of applying the graphical views extraction process is important for wider application of the proposed method.
@@
 In addition, we give the ratio of generated results that can pass the graphical views extraction process, which is denoted by Extraction Rate.
-The Pass@1 and the Extraction Rate of the generated results passing the graphical extraction process are given in Table 4 and Table 5.
-Table 4: The extraction rate of the generated results passing the graphical extraction process (C++).
+The Pass@1 and the Extraction Rate of the generated results passing the graphical extraction process are given in Table 5.
+Table 5: The extraction rate of the generated results passing the graphical extraction process.
 Generated Codes Pass@1 Extraction Rate
-Code-RAG 62.20 92.07
-CodeGRAG 64.02 92.68
-Table 5: The extraction rate of the generated results passing the graphical extraction process (Python).
-Generated Codes Pass@1 Extraction Rate
-Code-RAG 71.95 91.46
-CodeGRAG 77.44 96.95
+(C++) Code-RAG 62.20 92.07
+(C++) CodeGRAG 64.02 92.68
+(Python) Code-RAG 71.95 91.46
+(Python) CodeGRAG 77.44 96.95
 From the results, we could see that the extraction rates are high for codes to pass the graphical views extraction process, even under the situation where the Pass@1 ratios of the generated results are low.
 This indicates that the application range of the proposed method is wide.
-In addition, as the code RAG also offers performance gains, one could use multiple views, including raw code blocks and graphical representations, as the retrieval knowledge base.
+In addition, as the code RAG also offers performance gains, one could use multiple views as the retrieval knowledge.
+4 Related Work
+LLMs for NL2Code.
+The evolution of the Natural Language to Code translation (NL2Code) task has been significantly influenced by the development of large language models (LLMs).
+Initially, general LLMs like GPT-J (Radford et al., 2023), GPT-NeoX (Black et al., 2022), and LLaMA (Touvron et al., 2023a), despite not being specifically tailored for code generation, showed notable NL2Code capabilities due to their training on datasets containing extensive code data like the Pile (Gao et al., 2020) and ROOTS (Laurençon et al., 2022).
+To further enhance these capabilities, additional pretraining specifically focused on code has been employed.
+PaLM-Coder, an adaptation of the PaLM model (Chowdhery et al., 2023), underwent further training on an extra 7.8 billion code tokens, significantly improving its performance in code-related tasks.
+Similarly, Code LLaMA (Roziere et al., 2023) represents an advancement of LLaMA2 (Touvron et al., 2023b), benefiting from extended training on over 500 billion code tokens, leading to marked improvements over previous models in both code generation and understanding.
+These developments underscore the potential of adapting generalist LLMs to specific domains like NL2Code through targeted training, leading to more effective and efficient code translation solutions.
+Code Search.
+The code search methods can be summarized into three folds.
+Early methods utilizes sparse search to match the query and codes (Hill et al., 2011; Yang & Huang, 2017), which suffers from mismatched vocabulary due to the gap between natural language and codes.
+Neural methods (Cambronero et al., 2019; Gu et al., 2021) then focus on mapping the query and codes into a joint representation space for more accurate retrieval.
+With the success of pretrained language models, many methods propose to use pretraining tasks to improve the code understanding abilities and align different language spaces.
+For example, CodeBERT (Feng et al., 2020) is pretrained on NL-PL pairs of 6 programming languages with the masked language modeling and replaced token detection task.
+CodeT5 (Wang et al., 2021) supports both code-related understanding and generation tasks through bimodal dual generation.
+UniXcoder (Guo et al., 2022) integrates the aforementioned pretraining tasks, which is a unified cross-modal pre-trained model.
+As retrieval augmented generation (RAG) shows its significance in promoting the quality of LLMs generation, works in code RAG start to accumulate.
+Nashid et al.
+(2023); Lu et al.
+(2022) utilize the code blocks as the retrieved knowledge to inform the LLMs with similar code blocks for enhancement.
+Zhou et al.
+(2022) uses the programming related document to serve as the retrieval content, injecting auxiliary external programming knowledge into the LLMs generation.
+Code Representation.
+Early methods regard code snippets as sequences of tokens, assuming the adjacent tokens will have strong correlations.
+This line of methods (Harer et al., 2018; Ben-Nun et al., 2018; Feng et al., 2020; Ciniselli et al., 2021) take programming languages as the same with the natural language, using language models to encode the code snippets too.
+However, this ignoring of the inherent structure of codes leads to a loss of expressiveness.
+Methods that take the structural information of codes into consideration then emerge.
+Mou et al.
+(2016) used convolution networks over the abstract syntax tree (AST) extracted from codes.
+Alon et al.
+(2019) encoded paths sampled from the AST to represent codes.
+Further exploration into the graphical representation of codes (Allamanis et al., 2017) is conducted to better encode the structures of codes, where more intermediate states of the codes are considered.
 5 Conclusion
@@
 By integrating external structural knowledge, CodeGRAG enhances LLMs’ comprehension of code syntax and empowers them to generate complex and multi-lingual code with improved accuracy and fluency.
-The method’s demonstrated performance gains signal its potential to significantly elevate LLMs’ code generation abilities, including facilitating cross-lingual programming language generation such as C++ to Python.
-Overall, the findings warrant optimism regarding the transformative impact of CodeGRAG in advancing LLMs’ proficiency in code generation across various programming languages.
-Limitations
-In this paper, we propose a tuning-free method that can offer enhanced single round and cross-lingual code generation.
-Despite the efficiency and effectiveness, there are also limitations within this work.
-Firstly, dependency on the quality of the external knowledge base could be a potential concern.
-The quality of the external knowledge base could be improved with regular expression extraction on the noisy texts and codes.
-Secondly, the potential of combining the RAG with finetuning for improved generation is not explored in this work.
-Finetuning the LLMs with the assistance of CodeGRAG could be a future exploration.
-Appendix A Example of the inserted graphical view
-An illustration of the inserted graphical view is given below.
-{mdframed}
-Graph(
-num_nodes=’node’: 24,
-num_edges=(’node’, ’-0’, ’node’): 1, (’node’, ’-1’, ’node’): 1, (’node’, ’ArraySubscriptExpredge0’, ’node’): 1, (’node’, ’ArraySubscriptExpredge1’, ’node’): 1, (’node’, ’CXXOperatorCallExpredge1’, ’node’): 1, (’node’, ’CXXOperatorCallExpredge2’, ’node’): 2, (’node’, ’ImplicitCastExpredge0’, ’node’): 1, (’node’, ’UserDefineFun’, ’node’): 1, (’node’, ’falseNext’, ’node’): 1, (’node’, ’next’, ’node’): 5, (’node’, ’read’, ’node’): 10, (’node’, ’trueNext’, ’node’): 1, (’node’, ’write’, ’node’): 9,
-metagraph=[(’node’, ’node’, ’-0’), (’node’, ’node’, ’-1’), (’node’, ’node’, ’ArraySubscriptExpredge0’), (’node’, ’node’, ’ArraySubscriptExpredge1’), (’node’, ’node’, ’CXXOperatorCallExpredge1’), (’node’, ’node’, ’CXXOperatorCallExpredge2’), (’node’, ’node’, ’ImplicitCastExpredge0’), (’node’, ’node’, ’UserDefineFun’), (’node’, ’node’, ’falseNext’), (’node’, ’node’, ’next’), (’node’, ’node’, ’read’), (’node’, ’node’, ’trueNext’), (’node’, ’node’, ’write’)]
+Author Contributions
+If you’d like to, you may include a section for author contributions as is done in many journals. This is optional and at the discretion of the authors.
+Acknowledgments
+Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper.
+Appendix A Appendix
+You may include other additional sections here.

### v3

@@
 To take the extracted structural knowledge into the foundation models, we propose 1) a hard meta-graph prompt template to transform the challenging graphical representation into informative knowledge for tuning-free models and 2) a soft prompting technique that injects the domain knowledge of programming languages into the model parameters via finetuning the models with the help of a pretrained GNN expert model.
-Various experiments and ablations are done on four datasets including both the C++ and python languages to validate the hard meta-graph prompt, the soft prompting technique, and the effectiveness of the objectives for pretrained GNN expert.
-CodeGRAG improves the code generation ability of LLMs and can even offer performance gain for cross-lingual code generation.
-The implementation is available at https://anonymous.4open.science/r/Code-5970/ .
+CodeGRAG significantly improves the code generation ability of LLMs and can even offer performance gain for cross-lingual code generation.
+Implementation is available at https://anonymous.4open.science/r/Code-5970/ .
+CodeGRAG: Bridging the Gap between Natural Language and Programming Language via Graphical Retrieval Augmented Generation
+Kounianhua Du1, Jizheng Chen1, Renting Rui1, Huacan Chai1, Lingyue Fu1, Wei Xia2, Yasheng Wang2, Ruiming Tang2, Yong Yu1, Weinan Zhang1 1Shanghai Jiao Tong University, 2 Huawei Noah’s Ark Lab Shanghai, China {kounianhuadu, wnzhang}@sjtu.edu.cn
 1 Introduction
@@
 Each node in the tree represents a syntactic construct in the source code, such as a statement, an expression, or a declaration.
-ASTs are used in a variety of compiler construction and program analysis tasks, including: parsing, type checking, optimization, and program analysis.
 ASTs have good compactness and can represent the structure of the source code in a clear and concise way.
@@
 As for the node names, the middle figure displays the concrete types of nodes (operands) and the right figure displays the properties of nodes.
-An illustration of the composed graphical view is given in Figure 3.
-After obtaining the composed syntax graphs for code blocks in the retrieval pool, we use them to inform the general-purpose LLMs to bridge the gap between NL and PLs, where both the semantic level and the logic level information are preserved.
+An illustration of the composed graphical view is in Figure 3.
+After obtaining the composed syntax graphs, we use them to inform the general-purpose LLMs to bridge the gap between NL and PLs, where both the semantic level and the logic level information are preserved.
 2.3 Knowledge Querying
-Given a target problem to be completed, we generate informative query of it and use it to retrieve graphical knowledge from the constructed knowledge base. The process can be formulated as:
-$q\xleftarrow{}\text{QueryExtractor}(p),$ (3)
-[…]
-where $q$ denotes the query content, $p$ denotes the target problem, and $i$ is the returned index of the Top-1 relevant content stored in the constructed knowledge base.
-The main problems of the retrieval lie in: 1) how to design the informative query content and 2) how to align the different modalities.
-2.3.1 Query Extractor
-Since the styles of different code problems can diversify, the query content of the retrieval process matters.
-We consider the following contents: 1) Problem description, which describes the task to be completed by the target function code.
-Potential ambiguity and style diversity may exist among different problems set, which lead to a decrease in retrieval accuracy.
-2) Function declaration, which gives the function name and the input variables.
-Before knowledge querying, we first extract the problem description of each task to reduce the ambiguity and then concatenate it with the function declaration to serve as the query content, where the functionality and input format of the expected code block are contained.
-2.3.2 Retriever
+Given a target problem to be completed, we generate informative query of it and use it to retrieve graphical knowledge from the constructed knowledge base.
+We extract the problem description of each task to reduce the ambiguity and then concatenate it with the function declaration to serve as the query content, where the functionality and input format of the expected code block are contained.
 The query of the retrieval includes problem description $Q_{p}$ and function description $Q_{c}$, while each content of the retrieval pool includes raw code block $V_{c}$ and its graphical view $V_{g}$.
@@
 Therefore, we propose to use a meta-graph template to abstract the information of the graphical view.
-The abstracted meta-graph consists of the canonical edge types and node types, which describes the basic topology of the graphical view (Sun & Han, 2013), with the textual features obtained from the ASTs contained in the node and edge features.
+The abstracted meta-graph consists of the canonical edge types and node types, which describes the basic topology of the graphical view (Sun and Han, 2013), with the textual features obtained from the ASTs contained in the node and edge features.
 Then we use the meta-graph template to transform the retrieved graphical view into digestable knowledge and insert it into the final prompt for generation.
-As illustrated in Figure 5, the final prompt consists of three components: the system prompt illustrated in the blue part, the retrieved knowledge and hints illustrated in the green part, and the problem (including task description, function declaration, etc.) illustrated in the yellow part.
+As illustrated in Figure 5 in the Appendix, the final prompt consists of three components: the system prompt illustrated in the blue part, the retrieved knowledge and hints illustrated in the green part, and the problem (including task description, function declaration, etc.) illustrated in the yellow part.
 The three parts are concatenated to be fed into LLMs for knowledge augmented generation.
-Figure 4: Hard meta-graph prompt.
-Figure 5: Soft prompting.
 2.4.2 Soft Prompting with the Expert
@@
 To help the backbone LLMs to digest the graphical views, we tune the LLMs with the expert soft signal using supervised finetuning.
-The prompt for finetuning consists of the system prompt, retrieved knowledge where the expert encoded graphical view is contained using a token embedding, and task prompt, which is illustrated in Figure 5.
-Inference. After the finetuning stage, we used the tuned models to generate codes using the soft prompting template as illustrated in Figure 5.
+The prompt for finetuning consists of the system prompt, retrieved knowledge where the expert encoded graphical view is contained using a token embedding, and task prompt, which is illustrated in Figure 5 in the Appendix.
+Inference. After the finetuning stage, we used the tuned models to generate codes using the soft prompting template as illustrated in Figure 5 in the Appendix.
 3 Experiments
@@
     How is the compatibility of the graphical view?
-3.1 Setup
-In this paper, we evaluate CodeGRAG with the widely used HumanEval-X (Zheng et al., 2023) dataset, which is a multi-lingual code benchmark and CodeForce dataset in which we collect real-world programming problems from codeforces [^1] https://codeforces.com/ website.
-For CodeForce dataset we include problems categorized by different difficulty levels corresponding to the website and select $469$ problems of difficulty level A for testing.
-We use greedy decoding strategy to do the generation.
-The evaluation metric is Pass@1.
-For the size of retrieval pool, we use 11,913 C++ code snippets and 2,359 python code snippets.
-Due to the limited access, we do not use a large retrieval corpus for our experiment, which can be enlarged by other people for better performance.
-We also attach the graph extraction codes for both languages and all other expeirment codes here: https://anonymous.4open.science/r/Code-5970/
-For the fintuning details, the learning rate and weight decay for the expert GNN training is 0.001 and 1e-5, repectively.
-We apply 8-bit quantization and use LoRA for parameter-efficient fine-tuning.
-The rank of the low-rank matrices in LoRA is uniformly set to 8, alpha set to 16, and dropout is set to 0.05.
-The LoRA modules are uniformly applied to the Q and V parameter matrices of the attention modules in each layer of the LLM.
-All the three models are optimized using the AdamW optimizer.
-For the CodeContest dataset, totally 10609 datapoints are used, and for APPS dataset, 8691 data samples are used to train the model.
-3.2 Main Results
-The main results are summarized in Table 1 and Table 2.
 Table 1: Results of Hard Meta-Graph Prompt on Humaneval-X. (Pass@1)
@@
 (Multi-Lingual) Meta-Graph 64.02 77.44
-GPT-4omini N/A 65.24 78.66
+GPT-4omini N/A 63.41 78.66
 Code Block (Nashid et al., 2023; Lu et al., 2022) 65.24 78.66
@@
 Soft Prompting 11.09 30.26
-From the results, we can draw the following conclusions:
+3.1 Setup
+In this paper, we evaluate CodeGRAG with the widely used HumanEval-X (Zheng et al., 2023) dataset, which is a multi-lingual code benchmark and CodeForce dataset in which we collect real-world programming problems from codeforces [^1] https://codeforces.com/ website.
+For CodeForce dataset we include problems categorized by different difficulty levels corresponding to the website and select $469$ problems of difficulty level A for testing.
+We use greedy decoding strategy to do the generation.
+The evaluation metric is Pass@1.
+More details of the retrieval pool and the finetuning setting can be found in Section A in the Appendix.
+3.2 Main Results
+The main results are summarized in Table 1 and Table 2. From the results, we can draw the following conclusions.
 RQ1.
@@
 3.3 Impacts of the pretraining objectives for the expert GNN (RQ4)
-To study the effectiveness of the proposed pretraining objectives for the expert GNN, we remove each objective to yield different expert GNNs. The results are in Table 3.
 Table 3: Ablation studies on the GNN pretraining losses.
@@
 w/o Structure-Preserving 10.66 26.59
+To study the effectiveness of the proposed pretraining objectives for the expert GNN, we remove each objective to yield different expert GNNs. The results are in Table 3.
 From the results, we could see that both the Alignment and the Structure Preserving contribute to the expressiveness of the expert GNN model.
@@
 The code search methods can be summarized into three folds.
-Early methods utilizes sparse search to match the query and codes (Hill et al., 2011; Yang & Huang, 2017), which suffers from mismatched vocabulary due to the gap between natural language and codes.
+Early methods utilizes sparse search to match the query and codes (Hill et al., 2011; Yang and Huang, 2017), which suffers from mismatched vocabulary due to the gap between natural language and codes.
 Neural methods (Cambronero et al., 2019; Gu et al., 2021) then focus on mapping the query and codes into a joint representation space for more accurate retrieval.
@@
 5 Conclusion
-Despite the expanding role of LLMs in code generation, there are inherent challenges pertaining to their understanding of code syntax and their multi-lingual code generation capabilities.
-This paper introduces the Syntax Graph Retrieval Augmented Code Generation (CodeGRAG) to enhance LLMs for single round and cross-lingual code generation.
-CodeGRAG extracts and summarizes data flow and control flow information from codes, effectively bridging the gap between programming language and natural language.
-By integrating external structural knowledge, CodeGRAG enhances LLMs’ comprehension of code syntax and empowers them to generate complex and multi-lingual code with improved accuracy and fluency.
-Author Contributions
-If you’d like to, you may include a section for author contributions as is done in many journals. This is optional and at the discretion of the authors.
+Despite the expanding role of LLMs in code generation, there are inherent challenges pertaining to their understanding of code syntax.
+General large language models trained mainly on sequential-based natural language cannot well understand the structural-based programming language, e.g., the branching and jumping in codes.
+This paper proposes an effective way to build a graphical view of codes to better inform LLMs for code generation.
+To take the challenging structural graphical knowledge into LLMs, a meta-graph prompt is proposed for tuning-free models and a soft-prompting technique is proposed to inject the structural programming domain knowledge into the parameters of LLMs.
+By integrating external structural knowledge, CodeGRAG enhances LLMs’ comprehension of code syntax and empowers them to generate code with improved accuracy and fluency.
+Limitations
+In this paper, we propose a graphical retrieval augmented generation method that can offer enhanced code generation.
+Despite the efficiency and effectiveness, there are also limitations within this work.
+For example, dependency on the quality of the external knowledge base could be a potential concern.
+The quality of the external knowledge base could be improved with regular expression extraction on the noisy texts and codes.
 Acknowledgments
-Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper.
-Appendix A Appendix
-You may include other additional sections here.
+This document has been adapted by Emily Allaway from the instructions for earlier ACL and NAACL proceedings, including those for NAACL 2024 by Steven Bethard, Ryan Cotterell and Rui Yan, ACL 2019 by Douwe Kiela and Ivan Vulić, NAACL 2019 by Stephanie Lukin and Alla Roskovskaya, ACL 2018 by Shay Cohen, Kevin Gimpel, and Wei Lu, NAACL 2018 by Margaret Mitchell and Stephanie Lukin, Bib suggestions for (NA)ACL 2017/2018 from Jason Eisner, ACL 2017 by Dan Gildea and Min-Yen Kan, NAACL 2017 by Margaret Mitchell, ACL 2012 by Maggie Li and Michael White, ACL 2010 by Jing-Shin Chang and Philipp Koehn, ACL 2008 by Johanna D.
+Moore, Simone Teufel, James Allan, and Sadaoki Furui, ACL 2005 by Hwee Tou Ng and Kemal Oflazer, ACL 2002 by Eugene Charniak and Dekang Lin, and earlier ACL and EACL formats written by several people, including John Chen, Henry S.
+Thompson and Donald Walker.
+Additional elements were taken from the formatting instructions of the International Joint Conference on Artificial Intelligence and the Conference on Computer Vision and Pattern Recognition.
+Appendix A Implementation Details
+For the size of retrieval pool, we use 11,913 C++ code snippets and 2,359 python code snippets.
+Due to the limited access, we do not use a large retrieval corpus for our experiment, which can be enlarged by other people for better performance.
+We also attach the graph extraction codes for both languages and all other expeirment codes here: https://anonymous.4open.science/r/Code-5970/
+For the fintuning details, the learning rate and weight decay for the expert GNN training is 0.001 and 1e-5, repectively.
+We apply 8-bit quantization and use LoRA for parameter-efficient fine-tuning.
+The rank of the low-rank matrices in LoRA is uniformly set to 8, alpha set to 16, and dropout is set to 0.05.
+The LoRA modules are uniformly applied to the Q and V parameter matrices of the attention modules in each layer of the LLM.
+All the three models are optimized using the AdamW optimizer.
+For the CodeContest dataset, totally 10609 datapoints are used, and for APPS dataset, 8691 data samples are used to train the model.
+Appendix B Prompt Template
+Figure 4: Hard meta-graph prompt.
+Figure 5: Soft prompting.

### v4

@@
 Utilizing large language models to generate codes has shown promising meaning in software development revolution.
-Despite the intelligence shown by the general large language models, their specificity in code generation can still be improved due to the syntactic gap and mismatched vocabulary existing among natural language and different programming languages.
-In this paper, we propose CodeGRAG, a Graphical Retrieval Augmented Code Generation framework to enhance the performance of LLMs.
-CodeGRAG builds the graphical view of code blocks based on the control flow and data flow of them to fill the gap between programming languages and natural language, which can facilitate natural language based LLMs for better understanding of code syntax and serve as a bridge among different programming languages.
-To take the extracted structural knowledge into the foundation models, we propose 1) a hard meta-graph prompt template to transform the challenging graphical representation into informative knowledge for tuning-free models and 2) a soft prompting technique that injects the domain knowledge of programming languages into the model parameters via finetuning the models with the help of a pretrained GNN expert model.
+Despite the intelligence shown by the large language models, their specificity in code generation can still be improved due to the syntactic gap and mismatched vocabulary existing between natural language and programming languages.
+In this paper, we propose CodeGRAG, a Graphical Retrieval Augmented Code Generation framework that bridges the gap between NL and PL to enhance the performance of LLMs.
+CodeGRAG builds the graphical view of code blocks based on the control flow and data flow of them to better interpret the programming domain knowledge, which can facilitate natural language based LLMs for better understanding of code syntax and serve as a bridge among different programming languages.
+To take the extracted structural knowledge into the foundation models, we propose 1) a hard meta-graph prompt template to transform the challenging syntax graph into informative graphical view for tuning-free models and 2) a soft prompting technique that injects the domain knowledge of programming languages into model parameters via finetuning the models with the soft signals encoded by GNN expert model.
+Specifically, two constraints are designed to improve the alignment and structure expressiveness, contributing to the informativeness of the single-token-sized external <GraphEmb> for enhanced code generation.
 CodeGRAG significantly improves the code generation ability of LLMs and can even offer performance gain for cross-lingual code generation.
 Implementation is available at https://anonymous.4open.science/r/Code-5970/ .
-CodeGRAG: Bridging the Gap between Natural Language and Programming Language via Graphical Retrieval Augmented Generation
-Kounianhua Du1, Jizheng Chen1, Renting Rui1, Huacan Chai1, Lingyue Fu1, Wei Xia2, Yasheng Wang2, Ruiming Tang2, Yong Yu1, Weinan Zhang1 1Shanghai Jiao Tong University, 2 Huawei Noah’s Ark Lab Shanghai, China {kounianhuadu, wnzhang}@sjtu.edu.cn
 1 Introduction
-In recent years, large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023a) have shown great impact in various domains.
+In recent years, large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023) have shown great impact in various domains.
 Automated code generation emerges as a captivating frontier (Zheng et al., 2023; Roziere et al., 2023; Shen et al., 2023), promising to revolutionize software development by enabling machines to write and optimize code with minimal human intervention.
-However, syntactic gap and mismatched vocabulary among natural language (NL) and programming languages (PL) exist, hindering LLM’s performance on code generation.
-As illustrated in Figure 1, programming language (marked in blue) contains special tokens such as “int” or “++” that natural language (marked in yellow) doesn’t possess, leading to vocabulary mismatch.
-Besides, the relations between tokens in programming languages are often structural, e.g., the complex branching and jumps, whereas natural language is arranged simply in sequential manner, leading to syntactic gap.
-For example, in the control flow graph of the raw code (marked in pink), two “if” blocks (marked in purple) are adjacent and are executed sequentially under certain condition, but they appear to be intervaled in raw textual code.
-As discussed above, the innate structures of programming languages are different from that of the sequential-based natural language.
-The challenges of enhancing a general-purposed large language models for code-related tasks can be summarized into two folds.
+Code generation involves translating natural language instructions into structured and executable code in a target programming language.
+However, syntactic gap and mismatched vocabulary between natural language (NL) and programming languages (PL) exist, hindering LLM’s performance on code generation.
+As illustrated in Figure 1, programming language (marked in blue) contains special tokens such as “int” or “++” that natural language (marked in yellow) doesn’t possess with the same physical significance, leading to vocabulary mismatch.
+Besides, the relations among tokens in programming languages are often structural, e.g., the complex branching and jumps, whereas natural language is arranged simply in sequential manner, leading to syntactic gap.
+For example, in the control flow graph of the raw code (marked in pink), two “if” blocks (marked in purple) are adjacent and are executed sequentially under certain condition, while appearing distant in the raw textual code.
+As discussed above, the challenges of enhancing large language models for code-related tasks can be summarized into two folds.
+(C1) Effective representation of programming hints to interpret the inherent logic.
+Code, unlike natural language, possesses a well-defined structure that governs its syntax and semantics.
+This structure provides valuable information about the relationships between different parts of the code, the flow of execution, and the overall organization of the functions (Jiang et al., 2021; Guo et al., 2020).
+LLMs regard a code block as a sequence of tokens.
+By ignoring the inherent structure of codes, they miss out on essential cues that could help them better understand and generate code.
+Previous RAG methods (Lu et al., 2022; Su et al., 2024; Zhou et al., 2022) stimulate the code generation abilities of LLMs by using relevant code snippets or document as hints, failing to well represent the innate structures.
 Figure 1: Illustration of the gap between the programming language and the natural language.
-(C1) How to solve the gap between different languages and better interpret the inherent logic of code blocks.
-Code, unlike natural language, possesses a well-defined structure that governs its syntax and semantics.
-This structure provides valuable information about the relationships between different parts of the code, the flow of execution, and the overall organization of the functions Jiang et al.
-(2021); Guo et al.
-(2020).
-General-purpose LLMs regard a code block as a sequence of tokens.
-By ignoring the inherent structure of codes, they miss out on essential cues that could help them better understand and generate code.
-In addition, the multi-lingual code generation abilities of LLMs is challenging due to the gap among different programming languages.
-(C2) How to inject the innate knowledge of programming languages into general purpose large language models for enhancement.
-Despite the well representation of the programming knowledge, the ways to inject the knowledge into the NL-based foundation models is also challenging.
-The structural representation of codes could be hard to understand, which poses a challenge to the capability of the foundation models.
+(C2) Effective ways to digest the programming domain knowledge.
+In addition to the well representation of the programming knowledge, the ways to inject the knowledge into the NL-based LLMs is also challenging.
+The structural representation of codes and the unfamiliar vocabulary could be hard to understand, which poses a challenge to the capability of the foundation models.
 To solve the above challenges, we propose CodeGRAG, a graphical retrieval augmented generation framework for code generation.
-For (C1), we propose to interpret the code blocks using the composed graph based on the data-flow and control-flow of the code block, which extracts both the semantic level and the logical level information of the code.
-The composed graphical view could 1) better capture the innate structural knowledge of codes for NL-based language models to understand and 2) model the innate function of code blocks that bridging different programming languages.
-For (C2), we propose a meta-graph prompting technique for tuning-free models and a soft-prompting technique for tuned models.
-The meta-graph prompt summarizes the overall information of the extracted graphical view and transforms the challenging and noisy graphical representation into informative knowledge.
-The soft-prompting technique deals with the graphical view of codes with a pretrained GNN expert network and inject the processed knowledge embedding into the parameters of the general-purpose foundation models with the help of supervised finetuning.
+For (C1), we propose to interpret the code blocks using the graphical view composed from data-flow and control-flow of the code block to serve as the programming hints, which extracts both the semantic level and the logical level information of the code.
+The composed graphical view could 1) better capture the innate structural knowledge of codes for NL-based language models to understand and 2) model the innate function of code blocks that bridge different programming languages.
+For (C2), we propose a hard meta-graph prompting technique for tuning-free models and a soft-prompting technique for tunable models.
+The hard meta-graph prompt summarizes the overall information of the extracted graphical view and transforms the challenging and noisy graphical representation into informative knowledge.
+The soft-prompting technique injects the programming domain knowledge into LLMs by tuning them with soft signals from expert GNN-encoded graphical views, under the regularization of alignment and structure-preserving objectives.
+After the tuning, LLMs are capable to well digest the structural graphical views by taking in the soft <GraphEmb> at single-token cost, avoiding them from the vocabulary mismatch and syntactic gap problems.
 The main contributions of the paper can be summarized as follows:
-  • Novel GraphRAG framework for code generation.
-We propose CodeGRAG that bridges the gap among natural language and programming languages, transfers knowledge among different programming languages, and enhances the ability of LLMs for code generation.
-CodeGRAG requires only one calling of LLMs and can offer multi-lingual enhancement.
-  • Effective graphical view to inform and stimulate the structural programming knowledge of LLMs.
-We propose an effective graphical view to purify the semantic and logic knowledge from the code space, which offers more useful information than the raw code block and can summarize the cross-lingual knowledge.
-  • Effective soft prompting technique to preserve the programming domain knowledge and inject it into LLMs parameters.
-We propose an effective soft prompting technique, which injects the domain knowledge of programming languages into the model parameters via finetuning LLMs with the assistance of a pretrained GNN expert model.
+  • Identification of the vocabulary mismatch and syntactic gap problems in NL2Code process.
+We explain the identified problems with explicit examples and propose CodeGRAG, a novel graphical retrieval augmented generation framework, to tackle them for enhanced code generation.
+  • Effective graphical view to inform and stimulate the programming knowledge of LLMs.
+We propose an effective graphical view to serve as the programming hints, which offers more informative knowledge than the raw code block.
+  • Effective soft prompting technique to inject the programming domain knowledge into LLMs and help them better digest the programming hints.
+We propose a soft prompting technique, which injects the domain knowledge of programming languages into the model parameters via finetuning LLMs with the soft signals of GNN-encoded graphical representations.
+A two-hop contrastive objective for alignment along with a graph contrastive objective for structure preserving are designed to improve the expressiveness of the soft signal.
 2 Methodology
@@
 (Top) Knowledge Preparation.
-We extract the control flow and data flow of each external code block and compose them using the read-write signal to obtain the semantic and logical expression of each code block, which is then abstracted into graphical view as hard knowledge document and embedded into GraphEmb as soft knowledge document.
-The GraphEmb is encoded by a pretrained GNN expert model constrained by the alignment and structure preserving objectives.
-(Bottem) Retrieval Augmented Generation.
+We extract composed syntax graphs of external code blocks by composing the control flow and data flow of codes using the read-write signal, preserving the innate semantic and logical information.
+The composed graphs are then abstracted into graphical views as hard knowledge document and embedded into <GraphEmb>s as soft knowledge document.
+The <GraphEmb> is encoded by a pretrained GNN expert model constrained by the alignment and structure preserving objectives.
+(Bottom) Retrieval Augmented Generation.
 We extract query from the task input and retrieve from the external corpus.
 For tuning free models, we use the hard graphical view to stimulate the structural programming knowledge of LLMs for enhanced generation.
-For tunable models, we use the soft GraphEmb and inject the programming domain knowledge into LLMs parameters via finetuning them with the GNN expert signals.
+For tunable models, we use the soft <GraphEmb> and inject the programming domain knowledge into LLMs parameters via finetuning them with the GNN expert signals.
 The expert signals informed LLMs can then produce enhanced generation.
 2.1 Overview
-In this paper, we leverage both generative models and retrieval models to produce results that are both coherent and informed by the expert graphical knowledge of programming language.
-The overall process of CodeGRAG is illustrated in Figure 2, which mainly consists of three stages: graphical knowledge base preparation, knowledge querying, and graphical knowledge augmented generation.
+In this paper, we leverage both generative models and retrieval models to produce results that are coherent and informed by expert graphical knowledge of the programming language.
+The overall process of CodeGRAG is illustrated in Figure 2.
 2.2 Graphical Knowledge Base Preparation
-In this section, we discuss how to extract informative graphical views for code blocks.
-We analyze the syntax and control information of code blocks and extract their graphical views to better represent the codes.
-This process can be formulated as, $\forall c_{i}\in D_{\text{pool}}$:
-$g_{i}\xleftarrow{}\text{GraphExtractor}(c_{i}),$ (1)
-[…]
-where $c_{i}$ is the raw code block and $g_{i}$ is the corresponding extracted graphical view.
-To capture both the semantic and the logical information, we propose to combine the data flow graph (Aho et al., 2006) and the control flow graph (Allen, 1970) with the read-write signals (Long et al., 2022) to represent the code blocks, both of them are constructed on the base of the abstract syntax tree.
+To capture both the semantic and the logical information, we represent code blocks by combining the data flow graph (Aho et al., 2006) and the control flow graph (Allen, 1970) with the read-write signals (Long et al., 2022), both of them are constructed on the basis of the abstract syntax tree.
 Abstract Syntax Tree (AST).
@@
 An illustration of the composed graphical view is in Figure 3.
-After obtaining the composed syntax graphs, we use them to inform the general-purpose LLMs to bridge the gap between NL and PLs, where both the semantic level and the logic level information are preserved.
-2.3 Knowledge Querying
-Given a target problem to be completed, we generate informative query of it and use it to retrieve graphical knowledge from the constructed knowledge base.
-We extract the problem description of each task to reduce the ambiguity and then concatenate it with the function declaration to serve as the query content, where the functionality and input format of the expected code block are contained.
-The query of the retrieval includes problem description $Q_{p}$ and function description $Q_{c}$, while each content of the retrieval pool includes raw code block $V_{c}$ and its graphical view $V_{g}$.
-To expressively represent the components, we use the encoder $\phi(\cdot)$ of the pretrained NL2Code model to represent the problem description and code snippets. The retrieval function is:
-$\mathbf{h^{V}}$ $=\phi(V_{c}\|V_{g}),$ (3)
-$\mathbf{h^{Q}}$ $=\phi(Q_{p}\|Q_{c}),$ (4)
-Distance $=1-\frac{\mathbf{\mathbf{h^{Q}}\cdot\mathbf{h^{V}}}}{\|\mathbf{h^{Q}}\|\cdot\|\mathbf{h^{V}}\|}.$ (5)
-2.4 Graphical Knowledge Augmented Generation
-After we obtain the returned graphical view, we inject it to the foundation LLMs for graphical knowledge augmented generation.
-Since the graphical view is hard to understand, we propose 1) a meta-graph template to transform the graphical view into informative knowledge for tuning-free model and 2) a soft prompting technique to tune the foundation models for their better understanding of the graphical views with the assistance of an expert GNN model.
+After obtaining the composed syntax graphs, we use them to inform the LLMs to bridge the gap between NL and PLs, where both the semantic level and the logic level information are preserved.
+2.3 Graphical Retrieval Augmented Generation
+In this section, we discuss the ways to effectively utilize the constructed knowledge base.
+For the retrieval algorithm, we use CodeT5 to encode the query and corpus, measuring cosine distances among vectors to obtain the top-1 relevant document as the knowledge content.
+Since the focus of our paper is to discuss the representation of the used knowledge content but not the accuracy of ranking, we keep ranking algorithm the same for all used methods.
+For the representation of the retrieved knowledge content, since the composed syntax graph is hard to understand, we propose 1) a hard meta-graph template to transform the challenging knowledge into informative graphical view for tuning-free model and 2) a soft prompting technique to tune the foundation models for their better understanding of the graphical views with the assistance of an expert GNN model.
 2.3.1 Hard Meta-Graph Prompt
-The original graphical view of a code block could contain hundreds of nodes and edges.
-A full description of it could cost a overly long context, along with the understanding challenge posed by the long edge lists.
-Therefore, we propose to use a meta-graph template to abstract the information of the graphical view.
-The abstracted meta-graph consists of the canonical edge types and node types, which describes the basic topology of the graphical view (Sun and Han, 2013), with the textual features obtained from the ASTs contained in the node and edge features.
+The original composed syntax graph of a code block could contain hundreds of nodes and edges.
+A full description of it could cost an overly long context, along with the understanding challenge posed by the long edge lists.
+Therefore, we propose to use a meta-graph [^1] https://www.dgl.ai/dgl_docs/generated/dgl.DGLGraph.metagraph.html to abstract the original graph into informative graphical view.
+The abstracted meta-graph consists of the canonical edge types and node types, which describes the basic topology of the composed syntax graph (Sun and Han, 2013), with the textual features obtained from the ASTs contained in the node and edge features.
+The template for the meta-graph is displayed as below. {mdframed} Graph(num_nodes={$node\_type:\#nodes$},
+    num_edges=
+{$(src\_node\_type,edge\_type,$ $dst\_node\_type):\#edges$},
+[…]
 Then we use the meta-graph template to transform the retrieved graphical view into digestable knowledge and insert it into the final prompt for generation.
-As illustrated in Figure 5 in the Appendix, the final prompt consists of three components: the system prompt illustrated in the blue part, the retrieved knowledge and hints illustrated in the green part, and the problem (including task description, function declaration, etc.) illustrated in the yellow part.
+As illustrated in Figure 4 (a), the final prompt consists of three components: the system prompt illustrated in the blue part, the retrieved knowledge and hints illustrated in the green part, and the problem (including task description, function declaration, etc.) illustrated in the yellow part.
 The three parts are concatenated to be fed into LLMs for knowledge augmented generation.
+Figure 4: Prompt templates.
 2.3.2 Soft Prompting with the Expert
-Directly hard prompt to the LLMs poses a challenge to the digesting capability of the backbone LLMs, which could fail under the case where the backbone LLMs cannot well understand the graph components.
+Directly hard prompt to the LLMs poses a challenge to the digesting capability of the backbone LLMs, suffering from the vocabulary mismatch and syntactic gap problems.
 To compress the graphical knowledge into model parameters and help the backbone LLMs to better understand the programming language, we propose a soft prompting technique.
-The overall procedure can summarized into expert encoding of graphical views, finetuning with the expert signal, and inference.
-Expert Encoding of Graphical Views.
+The overall procedure can be summarized into expert encoding of graphical views, finetuning with the expert signal, and inference.
 We design a graph neural network to preserve the semantic and logical information of code blocks.
-The representation of each node $\mathbf{n}_{i}^{(0)}$ and edge $\mathbf{e}_{i}^{(0)}$ are first initialized with vectors corresponding to the node text and edge text encoded by $\phi_{1}$.
+The representation of each node $\mathbf{n}_{i}^{(0)}$ and edge $\mathbf{e}_{i}^{(0)}$ are first initialized with vectors corresponding to the node text and edge text encoded by $\phi_{1}$, which is a specialized coder model trained to align programming and natural languages.
 A message passing process is first conducted to fuse the semantic and structural information into each node representation.
@@
 $\mathbf{g}=\sum_{i}\text{softmax}(f_{\text{gate}}(\mathbf{n}_{i}^{L}))f_{\text{feat}}(\mathbf{n}_{i}^{L}).$ (6)
-The expert encoding network is optimized via the contrastive learning based self-supervised training, which includes the intra-modality contrastive learning and inter-modality contrastive learning.
-The intra-modality constrastive learning serves for preserving the modality information, while the inter-modality contrastive learning serves for modality alignment.
-  • Alignment Contrastive Learning.
-There are two types of alignment to be ensured: 1) NL-Code (NC) alignment and 2) Code-Graph (CG) alignment.
-We define the positive pairs for NC alignment purpose as $\mathcal{I}^{+}_{NC}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{i}^{Q}\rangle|i\in D_{\text{train}}\}$ and define the negative pairs for NC alignment purpose as $\mathcal{I}^{-}_{NC}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{j}^{Q}\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$
-    And we define the positive pairs for CG alignment purpose as $\mathcal{I}^{+}_{CG}=\{\langle\phi_{1}(c_{i}),\phi_{2}(g_{i})\rangle|i\in D_{\text{train}}\}$ and define the negative pairs for CG alignment purpose as $\mathcal{I}^{-}_{CG}=\{\langle\phi_{1}(c_{i}),\phi_{2}(g_{j})\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$
+The expert encoding network is optimized via the contrastive learning based self-supervised training.
+A two-hop contrastive learning objective is designed to align the code-NL-PL spaces, while a structure preserving contrastive learning is conducted to improve the expressiveness of the expert GNN and avoid the aligning process from being dominated by the NL side.
+  • Two-Hop Alignment Contrastive Learning. There are two types of alignment to be ensured: 1) NL-Code (NC) alignment and 2) Code-Graph (CG) alignment.
+    We define the positive pairs and negative pairs for NC alignment purpose as:
+    $\mathcal{I}^{+}_{NC}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{i}^{Q}\rangle$ $|i\in D_{\text{train}}\},$ (7)
+    $\mathcal{I}^{-}_{NC}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{j}^{Q}\rangle$ $|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$ (8)
+    And we define the positive pairs and negative pairs for CG alignment purpose as:
+    $\mathcal{I}^{+}_{CG}=\{\langle\phi_{1}(c_{i}),\phi_{2}(g_{i})\rangle|$ $i\in D_{\text{train}}\},$ (9)
+    $\mathcal{I}^{-}_{CG}=\{\langle\phi_{1}(c_{i}),\phi_{2}(g_{j})\rangle|$ $i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$ (10)
   • Structure Preserving Contrastive Learning.
@@
 Concretely, we corrupt each of the graphical view $g_{i}$ with the edge dropping operation to obtain its corrupted view $g_{i}^{\prime}$.
-The positive pairs for structure-preserving purpose are then designed as $\mathcal{I}^{+}_{\text{preserve}}=\{\langle\phi_{2}(g_{i}),\phi_{2}(g_{i}^{\prime})\rangle|i\in D_{\text{train}}\}$.
-The negative pairs for structure preserving purpose are designed as $\mathcal{I}^{-}_{\text{preserve}}=\{\langle\phi_{2}(g_{i}),\phi_{2}(g_{j}^{\prime})\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}$.
-Finetuning with the Expert Soft Signal.
-To help the backbone LLMs to digest the graphical views, we tune the LLMs with the expert soft signal using supervised finetuning.
-The prompt for finetuning consists of the system prompt, retrieved knowledge where the expert encoded graphical view is contained using a token embedding, and task prompt, which is illustrated in Figure 5 in the Appendix.
-Inference. After the finetuning stage, we used the tuned models to generate codes using the soft prompting template as illustrated in Figure 5 in the Appendix.
+The positive pairs and negative pairs for structure-preserving purpose are then designed as:
+    $\mathcal{I}^{+}_{\text{preserve}}$ =$=\{\langle\phi_{2}(g_{i}),\phi_{2}(g_{i}^{\prime})\rangle|i\in D_{\text{train}}\},$ (11)
+    $\mathcal{I}^{-}_{\text{preserve}}$ =$=\{\langle\phi_{2}(g_{i}),\phi_{2}(g_{j}^{\prime})\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$ (12)
+To help the backbone LLMs to digest the graphical views, we tune the LLMs with the expert soft signal.
+The prompt for finetuning consists of the system prompt, retrieved knowledge where the expert encoded graphical view is contained using a token embedding, and task prompt, as illustrated in Figure 4 (b).
+After the finetuning stage, we used the tuned models to generate codes using the soft prompting template.
 3 Experiments
@@
   • RQ3
-    Can soft prompting enhance the capability of the backbone LLMs? Does finetuning with the soft prompting outperforms the simple supervised finetuning?
+    Can soft prompting enhance the capability of the backbone LLMs? Does finetuning with the expert signals outperform the simple supervised finetuning ?
   • RQ4
@@
   • RQ6
-    How is the compatibility of the graphical view?
+    How is the compatibility of the graph extraction process?
+3.1 Setup
+In this paper, we evaluate CodeGRAG with the widely used HumanEval-X [^2] https://github.com/THUDM/CodeGeeX Zheng et al.
+(2023) dataset, competition-style CodeForce [^3] https://paperswithcode.com/sota/code-generation-on-apps and APPS [^4] https://arxiv.org/pdf/2105.09938v3 Hendrycks et al.
+(2021) datasets.
+We use greedy decoding strategy to do the generation.
+The evaluation metric is Pass@1.
+More details of the retrieval pool and the finetuning settings can be found in Section A.1 in the Appendix.
 Table 1: Results of Hard Meta-Graph Prompt on Humaneval-X. (Pass@1)
@@
 Soft Prompting 11.09 30.26
-3.1 Setup
-In this paper, we evaluate CodeGRAG with the widely used HumanEval-X (Zheng et al., 2023) dataset, which is a multi-lingual code benchmark and CodeForce dataset in which we collect real-world programming problems from codeforces [^1] https://codeforces.com/ website.
-For CodeForce dataset we include problems categorized by different difficulty levels corresponding to the website and select $469$ problems of difficulty level A for testing.
-We use greedy decoding strategy to do the generation.
-The evaluation metric is Pass@1.
-More details of the retrieval pool and the finetuning setting can be found in Section A in the Appendix.
 3.2 Main Results
@@
 From Table 2, we can see that finetuning with the expert soft prompting could offer more performance gain than that brought by simple supervised finetuning.
-This validates the effectiveness of the designed pretraining expert network and the technique of finetuning with soft prompting, which injects the programming domain knowledge into the LLMs parameters and inform the models with the structural information for gap filling.
-3.3 Impacts of the pretraining objectives for the expert GNN (RQ4)
-Table 3: Ablation studies on the GNN pretraining losses.
+This validates the effectiveness of the designed pretraining expert network and the technique of finetuning with soft prompting, which injects the programming domain knowledge into the LLMs parameters and informs the models with the structural information for gap filling.
+3.3 Impacts of the objectives for soft prompting (RQ4)
+Table 3: Ablation studies on the two objectives.
 Model Finetune CodeForce (C++) APPS (Python)
@@
 w/o Structure-Preserving 10.66 26.59
-To study the effectiveness of the proposed pretraining objectives for the expert GNN, we remove each objective to yield different expert GNNs. The results are in Table 3.
-From the results, we could see that both the Alignment and the Structure Preserving contribute to the expressiveness of the expert GNN model.
+Figure 5: T-sne visualization of soft signals trained with different objectives.
+To study the effectiveness of the proposed pretraining objectives for the expert GNN, we remove each objective to yield different expert GNNs and visualize the hidden representations of codes with T-SNE.
+The results are in Table 3 and Figure 5.
+From the results, we could see that both the Alignment and the Structure Preserving contribute to the expressiveness of the soft signal.
 The alignment pretraining objective helps to promote the alignment among natural language, programming language, and their graphical views.
@@
 The results are presented in Table 4.
-Table 4: The impacts of the graph components.
+Table 4: Impacts of the graph components.
 Datasets Python C++
@@
 From the results, we can observe that 1) the edge features matter the most in constructing the structural view of code blocks for enhancement, 2) the type of nodes expresses the most in representing operands information, and 3) the overall structure of the graphical view also gives additional information.
-3.5 Compatibility Discussion of the Graphical Views(RQ6)
+3.5 Compatibility Discussion (RQ6)
 Despite the effectiveness of the proposed graphical views to represent the code blocks, the flexibility and convenience of applying the graphical views extraction process is important for wider application of the proposed method.
 In this section, we discuss the compatibility of CodeGRAG.
-First of all, the extraction process of all the graphical views are front-end.
-Therefore, this extraction process applies to a wide range of code, even error code.
-One could also use convenient tools to reformulate the code and improve the pass rate of the extraction process.
-In addition, we give the ratio of generated results that can pass the graphical views extraction process, which is denoted by Extraction Rate.
-The Pass@1 and the Extraction Rate of the generated results passing the graphical extraction process are given in Table 5.
-Table 5: The extraction rate of the generated results passing the graphical extraction process.
-Generated Codes Pass@1 Extraction Rate
+Table 5: Graph extraction rate (ER) of the generated results.
+Generated Codes Pass@1 ER
 (C++) Code-RAG 62.20 92.07
@@
 (Python) CodeGRAG 77.44 96.95
+We give the ratio of generated results that can pass the graphical views extraction process, which is denoted by Extraction Rate.
+The Pass@1 and the Extraction Rate of the generated results passing the graphical extraction process are given in Table 5.
 From the results, we could see that the extraction rates are high for codes to pass the graphical views extraction process, even under the situation where the Pass@1 ratios of the generated results are low.
 This indicates that the application range of the proposed method is wide.
-In addition, as the code RAG also offers performance gains, one could use multiple views as the retrieval knowledge.
+In addition, the extraction process of all the graphical views are front-end.
+Therefore, this extraction process applies to a wide range of code, even error code.
+One could also use regularization tools to reformulate the code and improve the pass rate of the extraction process.
 4 Related Work
-LLMs for NL2Code.
-The evolution of the Natural Language to Code translation (NL2Code) task has been significantly influenced by the development of large language models (LLMs).
-Initially, general LLMs like GPT-J (Radford et al., 2023), GPT-NeoX (Black et al., 2022), and LLaMA (Touvron et al., 2023a), despite not being specifically tailored for code generation, showed notable NL2Code capabilities due to their training on datasets containing extensive code data like the Pile (Gao et al., 2020) and ROOTS (Laurençon et al., 2022).
-To further enhance these capabilities, additional pretraining specifically focused on code has been employed.
-PaLM-Coder, an adaptation of the PaLM model (Chowdhery et al., 2023), underwent further training on an extra 7.8 billion code tokens, significantly improving its performance in code-related tasks.
-Similarly, Code LLaMA (Roziere et al., 2023) represents an advancement of LLaMA2 (Touvron et al., 2023b), benefiting from extended training on over 500 billion code tokens, leading to marked improvements over previous models in both code generation and understanding.
-These developments underscore the potential of adapting generalist LLMs to specific domains like NL2Code through targeted training, leading to more effective and efficient code translation solutions.
 Code Search.
 The code search methods can be summarized into three folds.
-Early methods utilizes sparse search to match the query and codes (Hill et al., 2011; Yang and Huang, 2017), which suffers from mismatched vocabulary due to the gap between natural language and codes.
+Early methods utilize sparse search to match the query and codes (Hill et al., 2011; Yang and Huang, 2017), which suffers from mismatched vocabulary due to the gap between NL and PL.
 Neural methods (Cambronero et al., 2019; Gu et al., 2021) then focus on mapping the query and codes into a joint representation space for more accurate retrieval.
@@
 Further exploration into the graphical representation of codes (Allamanis et al., 2017) is conducted to better encode the structures of codes, where more intermediate states of the codes are considered.
-5 Conclusion
-Despite the expanding role of LLMs in code generation, there are inherent challenges pertaining to their understanding of code syntax.
-General large language models trained mainly on sequential-based natural language cannot well understand the structural-based programming language, e.g., the branching and jumping in codes.
-This paper proposes an effective way to build a graphical view of codes to better inform LLMs for code generation.
-To take the challenging structural graphical knowledge into LLMs, a meta-graph prompt is proposed for tuning-free models and a soft-prompting technique is proposed to inject the structural programming domain knowledge into the parameters of LLMs.
-By integrating external structural knowledge, CodeGRAG enhances LLMs’ comprehension of code syntax and empowers them to generate code with improved accuracy and fluency.
-Limitations
+5 Limitations
 In this paper, we propose a graphical retrieval augmented generation method that can offer enhanced code generation.
@@
 The quality of the external knowledge base could be improved with regular expression extraction on the noisy texts and codes.
-Acknowledgments
-This document has been adapted by Emily Allaway from the instructions for earlier ACL and NAACL proceedings, including those for NAACL 2024 by Steven Bethard, Ryan Cotterell and Rui Yan, ACL 2019 by Douwe Kiela and Ivan Vulić, NAACL 2019 by Stephanie Lukin and Alla Roskovskaya, ACL 2018 by Shay Cohen, Kevin Gimpel, and Wei Lu, NAACL 2018 by Margaret Mitchell and Stephanie Lukin, Bib suggestions for (NA)ACL 2017/2018 from Jason Eisner, ACL 2017 by Dan Gildea and Min-Yen Kan, NAACL 2017 by Margaret Mitchell, ACL 2012 by Maggie Li and Michael White, ACL 2010 by Jing-Shin Chang and Philipp Koehn, ACL 2008 by Johanna D.
-Moore, Simone Teufel, James Allan, and Sadaoki Furui, ACL 2005 by Hwee Tou Ng and Kemal Oflazer, ACL 2002 by Eugene Charniak and Dekang Lin, and earlier ACL and EACL formats written by several people, including John Chen, Henry S.
-Thompson and Donald Walker.
-Additional elements were taken from the formatting instructions of the International Joint Conference on Artificial Intelligence and the Conference on Computer Vision and Pattern Recognition.
-Appendix A Implementation Details
-For the size of retrieval pool, we use 11,913 C++ code snippets and 2,359 python code snippets.
+6 Conclusion
+In this paper, we locate the syntactic gap and vocabulary mismatch problems in NL2Code process and propose CodeGRAG to alleviate the above problems for enhanced code generation.CodeGRAG effectively represent the programming hints to inform LLMs for code generation and improve their understanding of programming domain knowledge.
+Concretely, we propose to extract the semantic and logic knowledge of codes to serve as the programming hints and design proper ways to take in the knowledge: 1) a meta-graph prompt for tuning-free models and 2) a soft-prompting technique to inject the structural programming domain knowledge into the parameters of LLMs for tunable models.
+Two objectives are designed to improve the alignment and structural expressiveness of the soft signals, contributing to the informativeness of the single-token-sized <GraphEmb> for enhanced code generation.
+By integrating properly-represented external knowledge, CodeGRAG enhances LLMs’ comprehension of codes and empowers them to generate code with improved accuracy.
+Acknowledgments and Disclosure of Funding
+Use unnumbered first level headings for the acknowledgments.
+All acknowledgments go at the end of the paper before the list of references.
+Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work).
+More information about this disclosure can be found at: https://neurips.cc/Conferences/2024/PaperInformation/FundingDisclosure.
+Do not include this section in the anonymized submission, only in the final paper.
+You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
+Appendix A Appendix
+A.1 Implementation Details
+For the size of the retrieval pool, we use $11,913$ C++ code snippets and $2,359$ Python code snippets.
 Due to the limited access, we do not use a large retrieval corpus for our experiment, which can be enlarged by other people for better performance.
-We also attach the graph extraction codes for both languages and all other expeirment codes here: https://anonymous.4open.science/r/Code-5970/
-For the fintuning details, the learning rate and weight decay for the expert GNN training is 0.001 and 1e-5, repectively.
+We also attach the graph extraction codes for both languages and all other experiment codes here: https://anonymous.4open.science/r/Code-5970/
+For the finetuning details, the learning rate and weight decay for the expert GNN training are 0.001 and 1e-5, respectively.
 We apply 8-bit quantization and use LoRA for parameter-efficient fine-tuning.
@@
 The LoRA modules are uniformly applied to the Q and V parameter matrices of the attention modules in each layer of the LLM.
-All the three models are optimized using the AdamW optimizer.
-For the CodeContest dataset, totally 10609 datapoints are used, and for APPS dataset, 8691 data samples are used to train the model.
-Appendix B Prompt Template
-Figure 4: Hard meta-graph prompt.
-Figure 5: Soft prompting.
+All three models are optimized using the AdamW optimizer.
+For the CodeContest dataset, total $10,609$ data points are used, and for the APPS dataset, $8,691$ data samples are used to train the model.
+NeurIPS Paper Checklist
+The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact.
+Do not remove the checklist: The papers not including the checklist will be desk rejected.
+The checklist should follow the references and follow the (optional) supplemental material.
+The checklist does NOT count towards the page limit.
+Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
+  • You should answer [Yes] , [No] , or [N/A] .
+  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
+  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
+The checklist answers are an integral part of your paper submission.
+They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers.
+You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
+The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation.
+While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used").
+In general, answering "[No] " or "[N/A] " is not grounds for rejection.
+While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate.
+All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix.
+If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
+IMPORTANT, please:
+  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
+  • Keep the checklist subsection headings, questions/answers and guidelines below.
+  • Do not modify the questions and only use the provided macros for your answers.
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: See Section 3.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: See Section 5.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: N/A.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: The implementation is available at https://anonymous.4open.science/r/Code-5970/. However, we do not provide the API needed to reproduce the results.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: The implementation is available at https://anonymous.4open.science/r/Code-5970/.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: See Section 3.1 and Appendix A.1.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34. Justification: N/A.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: See Appendix A.1.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: The research conforms with the Neurips Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: See Section 6.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: N/A.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: The existing assets used are properly cited.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: Codes are public.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: N/A.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: N/A.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.